### v1

# An explicit form of Ingham’s zero density estimate

Abstract.
Ingham (1940) proved that $N(\sigma,T)\ll T^{3(1-\sigma)/(2-\sigma)}\log^{5}{T}$, where $N(\sigma,T)$ counts the number of the non-trivial zeros $\rho$ of the Riemann zeta-function with $\Re\{\rho\}\geq\sigma\geq 1/2$ and $0<\Im\{\rho\}\leq T$. We provide an explicit version of this result with the exponent $3$ of the logarithmic factor. In addition, we also provide an explicit estimate with asymptotically correct main term for the fourth power moment of the Riemann zeta-function on the critical line.
Key words and phrases:
Riemann zeta-function, Zero density estimates, Moments of the Riemann zeta-function
2020 Mathematics Subject Classification:
11M06, 11M26
1. Introduction
Let $\zeta(s)$ be the Riemann zeta-function and let $\rho=\beta+\mathrm{i}\gamma$ be its non-trivial zero, that is, $\zeta(\rho)=0$ with $0\leq\beta\leq 1$ and $\gamma\in\mathbb{R}$. Let $N(T)$ be the number (counted with multiplicities) of all non-trivial zeros with $0<\gamma\leq T$, while for $\sigma\geq 0$ and $T>0$ define $N(\sigma,T)\mathrel{\mathop{:}}=\#\left\{\rho\colon\zeta(\rho)=0,\beta\geq\sigma,0<\gamma\leq T\right\}$, where the zeros are again counted with multiplicities. Trivially, the function $N(\sigma,T)$ is decreasing in $\sigma$ and increasing in $T$, $N(\sigma,T)\leq N(T)\sim\frac{1}{2\pi}T\log{T}$ for $20\leq\sigma\leq 1/$ 2, $N(\sigma,T)\leq\frac{1}{2}N(T)$ for $11/2<\sigma<$ 1 and $N(\sigma,T)=0$ for $\sigma\geq 1$, while the Riemann Hypothesis (RH) asserts that $N(\sigma,T)=N(T)$ for $20\leq\sigma\leq 1/$ 2 and $N(\sigma,T)=0$ for $\sigma>1/2$. Since RH has been verified up to $H_{\mathrm{RH}}\mathrel{\mathop{:}}=3\cdot 10^{12}$ by Platt and Trudgian [PT21], we have $N(\sigma,T)=0$ for $\sigma>1/2$ and $T\leq H_{\mathrm{RH}}$. Consequently, we can restrict ourselves to the range of $1/2\leq\sigma\leq 1$ and $T\geq H_{\mathrm{RH}}$ throughout this paper.
Loosely speaking, a non-trivial estimate on $N(\sigma,T)$ is known as zero density estimate. Estimates of the form [^1] As usual, $A(\sigma)$ is defined as “infimum of all such exponents”, see [TTY25, Definition 37] for a precise definition.
$N(\sigma,T)\ll T^{A(\sigma)(1-\sigma)}\log^{B}{T}\quad\textrm{or}\quad N(\sigma,T)\ll_{\varepsilon}T^{A(\sigma)(1-\sigma)+\varepsilon}$ (1)
with $A(\sigma)\leq 1/(1-\sigma)$, $B\geq 0$, $\varepsilon>0$ and the implied constants being uniform in $\sigma$ and $T$, are of great interest in the distribution theory of prime numbers [Ivi03, Sta26]. The Lindelöf Hypothesis (LH), which is believed to be a weaker statement than RH, implies $A(\sigma)\leq 2$ in the expression on the right-hand side of (1) for all $\sigma\in[1/2,1]$, and this is known as the density hypothesis [Ivi03, p. 47]. Bourgain [Bou00] showed that the density hypothesis holds for $\sigma\geq 25/32=0.78125$, and substantial progress has been made in obtaining much better upper bounds for $A(\sigma)$ in this region, see [TTY25, Table 2 on p. 33] for the current best results. On the other hand, region $0.761/2\leq\sigma<19/25=$ 0.76 is dominated by two results, namely that
$A(\sigma)\leq\min\left\{\frac{3}{2-\sigma},\frac{15}{5\sigma+3}\right\},$
which is due to Ingham [Ing40] and Guth–Maynard [GM24], respectively. Note that $3/(2-\sigma)\leq 15/(5\sigma+3)$ for $\sigma\in[1/2,0.7]$. This means that, apart from Selberg’s zero density estimate [^2] His result is sharper than [Ing40] for $\sigma$ close to the critical line, e.g., when $0\leq\sigma-1/2\ll\log{\log{T}}/\log{T}$. [Sel46] (see also [Bal17, Sim20]), Huxley’s refinement [Hux72, Chapter 23]
$N(\sigma,T)\ll\left(\sigma-\frac{1}{2}\right)^{-\frac{1}{3}}T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{6}{2-\sigma}}$ (2)
of Ingham’s later [^3] This is to distinguish between $N(\sigma,T)\ll T^{3(1-\sigma)/(2-\sigma)}\log^{5}{T}$ and Ingham’s earlier zero density estimates, see [Ing37], such as $N(\sigma,T)\ll T^{8(1-\sigma)/3}\log^{5}{T}$. zero density estimate is currently the best known result on $N(\sigma,T)$ for $1/2<\sigma\leq 0.7$.
In order to obtain effective and unconditional results that are related to the distribution of prime numbers, explicit versions of (1) are often required, see [FKS23, JY23, CHJ25]. Recently, several relevant results have appeared in the literature, see [Kad13, KLN18, Bel25, Cho24] for example. The aim of this paper is to provide the first explicit version of (2), see Corollary 1, where we also made an improvement over the logarithmic factor.
Corollary 1.
Let $T\geq H_{\mathrm{RH}}$. Then
$N(\sigma,T)\leq\mathfrak{B}_{1}T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+\mathfrak{B}_{2}\log^{2}{T}+\mathfrak{B}_{3}\log{T}$ (3)
for $\sigma_{1}\leq\sigma\leq\sigma_{2}$, where the values for $\mathfrak{B}_{1}$, $\mathfrak{B}_{2}$ and $\mathfrak{B}_{3}$ are given by Table 1.
The intervals $[\sigma_{1},\sigma_{2}]$ in Table 1 are elements of the set $\left\{[n/32,(n+1)/32]\colon n\in\{16,\ldots,31\}\right\}$. This choice is justified by the fact that (3) is asymptotically better than $N(\sigma,T)\ll T^{8(1-\sigma)/3}\log^{5}{T}$ for $\sigma\in[1/2,0.875]$ and the latter zero density estimate [^4] We are comparing our result to [KLN18], where a variant of such a zero density estimate was proved. is asymptotically better than the trivial one for $\sigma\in(0.625,1)$. Corollary 1 implies that
$N(\sigma,T)\leq 8.604\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+9.461\log^{2}{T}+167.8\log{T},\quad\sigma\in[0.500,0.625],\;T\geq 3\cdot 10^{12},$
$N(\sigma,T)\leq 22.44\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+8.290\log^{2}{T}+147.0\log{T},\quad\sigma\in[0.625,0.875],\;T\geq 3\cdot 10^{12}.$
The above estimates improve [Cho24, Theorem 1.1 and Table 1] for $\sigma\in[0.5,0.667]$ and the latter one improves [KLN18, Theorem 1.1 and Table 1] for $\sigma\in[0.625,0.875]$, for $T\geq 3\cdot 10^{12}$.
$\sigma_{1}$ $\sigma_{2}$ $d$ $\kappa$ $\delta$ $\log_{10}A_{0}$ $\mathfrak{B}_{1}$ $\mathfrak{B}_{2}$ $\mathfrak{B}_{3}$
$0.50000$ $0.53125$ $0.28$ $0.36$ $0.012$ $10.3$ $5.360$ $9.461$ $167.8$
$0.53125$ $0.56250$ $0.28$ $0.34$ $0.025$ $10.4$ $6.380$ $9.168$ $162.6$
$0.56250$ $0.59375$ $0.28$ $0.32$ $0.038$ $10.4$ $7.453$ $8.875$ $157.4$
$0.59375$ $0.62500$ $0.28$ $0.30$ $0.051$ $10.4$ $8.604$ $8.582$ $152.2$
$0.62500$ $0.65625$ $0.28$ $0.28$ $0.066$ $10.4$ $9.848$ $8.290$ $147.0$
$0.65625$ $0.68750$ $0.28$ $0.26$ $0.080$ $10.4$ $11.20$ $7.997$ $141.9$
$0.68750$ $0.71875$ $0.28$ $0.24$ $0.096$ $10.6$ $12.65$ $7.704$ $136.7$
$0.71875$ $0.75000$ $0.28$ $0.22$ $0.112$ $10.6$ $14.23$ $7.411$ $131.5$
$0.75000$ $0.78125$ $0.28$ $0.20$ $0.129$ $10.6$ $15.95$ $7.118$ $126.3$
$0.78125$ $0.81250$ $0.28$ $0.18$ $0.147$ $10.6$ $17.84$ $6.826$ $121.1$
$0.81250$ $0.84375$ $0.28$ $0.16$ $0.166$ $10.6$ $19.97$ $6.533$ $115.9$
$0.84375$ $0.87500$ $0.28$ $0.15$ $0.187$ $10.6$ $22.44$ $6.240$ $110.7$
$0.87500$ $0.90625$ $0.28$ $0.13$ $0.209$ $10.8$ $25.51$ $5.947$ $105.5$
$0.90625$ $0.93750$ $0.29$ $0.11$ $0.234$ $10.8$ $29.71$ $5.463$ $96.87$
$0.93750$ $0.96875$ $0.29$ $0.09$ $0.261$ $10.8$ $36.09$ $5.180$ $91.86$
$0.96875$ $1.00000$ $0.29$ $0.07$ $0.292$ $10.8$ $46.92$ $4.897$ $86.84$
Table 1. The values for $\mathfrak{B}_{1}$, $\mathfrak{B}_{2}$ and $\mathfrak{B}_{3}$ from Corollary 1 for the corresponding $\sigma\in[\sigma_{1},\sigma_{2}]$, as well as the optimized values of several parameters from Theorem 1.
The precise result from which Corollary 1 follows is stated as Theorem 1 in Section 2.1. Its proof is based on Ingham’s original approach [Ing40], and relies [^5] Two other important ingredients are Gabriel’s convexity theorem (Theorem 4) and inequality (32) that is implied by Littlewood’s lemma. One can avoid them, usually for the cost of producing worse log-factor in (2), see [Ivi03, Chapter 11] and also [Mon71, Chapter 12] for generalisations to Dirichlet $L$-functions. In fact, Huxley established (2) for this class of $L$-functions. on an upper bound for the fourth power moment of $\zeta(s)$ on the critical line. Let $k$ be a positive real number and $0\leq A\leq B$, and define
$\mathcal{M}_{k}(A,B)\mathrel{\mathop{:}}=\int_{A}^{B}\left|\zeta\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2k}\mathrm{d}t.$ (4)
In particular, we set $\mathcal{M}_{k}(T)\mathrel{\mathop{:}}=\mathcal{M}_{k}(0,T)$ for $T\geq 0$. The LH trivially implies $\mathcal{M}_{k}(T)\ll_{\varepsilon}T^{1+\varepsilon}$, and if the latter is true for some strictly increasing and unbounded sequence of $k$’s, then LH holds [Tit86, Theorem 13.2]. However, much stronger statement
$\mathcal{M}_{k}(T)\asymp_{k}T\left(\log{T}\right)^{k^{2}}$ (5)
is believed to be true. Unconditionally, the upper bound in (5) is known to hold for $0<k\leq 2$, see [HRS19], while the lower bound is true for all $k\geq 1$, see [RS13]. On RH, the upper bound in (5) also holds for $k>2$, see [Har13], while the lower bound is also true for $10<k<$ 1, see [HB81]. The random matrix theory predicts the asymptotic form of (5) for $k\in\mathbb{N}$, that is,
$\mathcal{M}_{k}(T)=Tp_{k}(\log{T})+O_{\varepsilon}\left(T^{1/2+\varepsilon}\right)$ (6)
for $\varepsilon>0$, where $p_{k}(x)$ is a polynomial of degree $k^{2}$ whose coefficients can be effectively computed, see [CFK+05]. Polynomials $p_{k}(x)$ from (6) are known only for $k\in\{1,2\}$: $p_{1}(x)=x+2\gamma-1-\log{(2\pi)}$ where $\gamma$ is the Euler–Mascheroni constant is due to Ingham [Ing26], and $p_{2}(x)=\frac{1}{2\pi^{2}}x^{4}+\sum_{k=1}^{4}a_{4-k}x^{4-k}$ is due to Heath-Brown [HB79], where his [^6] His error term is of order $T^{7/8+\varepsilon}$. This was improved in [IM95], see also [PT25], to $T^{2/3+\varepsilon}$ by using methods from the spectral theory of the Riemann zeta-function. values for $a_{k}$, $k\in\{0,1,2,3\}$, coincide with those predicted by the random matrix theory. It should be mentioned that Ingham proved in the same paper [Ing26] that $p_{2}(x)=\frac{1}{2\pi^{2}}x^{4}+O(x^{3})$, i.e.,
$\mathcal{M}_{2}(T)=\frac{1}{2\pi^{2}}T\log^{4}{T}+O\left(T\log^{3}{T}\right),$ (7)
thus providing an asymptotic formula for $\mathcal{M}_{2}(T)$. Recently, explicit estimates for the second power moment $\mathcal{M}_{1}(T)$ appeared in [DZA22] and [SS23] that are of the form (6) but with better error terms. To the best of the authors knowledge, no explicit estimates for the fourth power moment $\mathcal{M}_{2}(T)$ have appeared in the literature. In this paper, we are providing them that are of the form $\ll T\log^{4}{T}$ and (7), see Corollary 2.
Corollary 2.
We have
$\mathcal{M}_{2}(T)\leq\frac{1}{\pi^{2}}T\log^{4}{\frac{T}{2}}+20.7225\cdot T\log^{\frac{7}{2}}{\frac{T}{2}}+1.9532\cdot 10^{6}$ (8)
for $T\geq 3\cdot 10^{3}$, and
$-\left(48.942+\frac{10^{5}\log{T}}{\pi^{2}T}\right)T\log^{3}{\frac{T}{2}}\leq\mathcal{M}_{2}(T)-\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}\leq 48.801\cdot T\log^{3}{\frac{T}{2}}+3.0592\cdot 10^{10}$ (9)
for $T\geq 10^{5}$.
The precise result from which Corollary 2 follows is stated as Corollary 4 in Section 2.2, which in turn is a further consequence of Theorem 2. Our proof of Theorem 2 follows Ramachandra’s proof [Ram75], see Section 5.1 for details. In addition to zero density estimates, results on the fourth power moment, as well as other higher power moments, are used in the Dirichlet divisor problem [Tit86, Section 12.3] and also to count integral ideals in abelian number fields [LLM25].
As an easy consequence of [DZA22] and Corollary 2 we also derive an explicit upper bound for the third power moment, see Corollary 3.
Corollary 3.
We have
$\mathcal{M}_{\frac{3}{2}}(T)\leq\left(\frac{1}{2\pi^{2}}+\frac{241.03}{\log{T_{0}}}+\frac{132.1}{\log^{2}{T_{0}}}\right)^{\frac{1}{2}}T\log^{\frac{5}{2}}{T}$
for $T\geq T_{0}\geq 10^{5}$.
The exponent of the logarithm is for $1/4$ larger than the conjectured value, see (5). It is interesting that this is still the best known unconditional result on $\mathcal{M}_{\frac{3}{2}}(T)$.
Although not pursued here, Corollary 1 can be generalized to Dirichlet $L$-functions by following the approach presented here; however, substantial work should be done on generalizing Lemmas 58 and 13. These and possible extensions to other $L$-functions are currently under investigation.
The outline of this paper is as follows. In Section 2 we state the main results, namely Theorem 1 for the zero density estimate and Theorem 2 for the fourth power moment estimate, and prove the above corollaries. In Section 3 we prove Gabriel’s convexity theorem, together with two applications (Corollaries 6 and 7), and in Section 4 we prove Theorem 1. Section 5 is dedicated to the proof of Theorem 2.
Acknowledgements
The authors thank Andrew Yang and Tim Trudgian for their interest in the work.
2. The main results
In this section, we are stating the precise results that are related to a zero density estimate (3), see Theorem 1, and to the fourth power moment estimates (8) and (9), see Theorem 2. In addition, proofs of Corollaries 12 and 3 are also provided.
2.1. The zero density estimate
The proof of Theorem 1 will be presented in Section 4.
Theorem 1.
Let $T\geq H\geq H_{0}\geq 1002$, $T\geq T_{0}\geq\max\left\{e^{2d},e^{e}\right\}$ for $d>0$, $0<\delta\leq\frac{1}{2}\log{(hT_{0})}$ for $h\geq 10^{9}/T_{0}$, and $\kappa>0$. Then
$N(\sigma,T)-N(\sigma,H)$ $\leq T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\frac{\mathfrak{A}_{1}}{2\pi d}\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}}\left(\log{T}\right)^{\frac{6-5\sigma}{2-\sigma}}+\frac{\mathfrak{A}_{2}}{2\pi d}\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}\log{T}\right)$
$+\left(\mathfrak{A}_{3}\cdot\log{(HT)}+\mathfrak{A}_{4}\cdot\log{(hT)}+\mathfrak{A}_{5}\right)\log{T}$
for $1/2+d/\log{T}<\sigma_{1}\leq\sigma\leq\sigma_{2}\leq 1$. Functions $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$ are defined by
$\mathfrak{A}_{1}$ $=\mathfrak{A}_{1}(\sigma_{1},\sigma_{2},T_{0},d,L_{0},L_{1},K_{1},K_{3})\mathrel{\mathop{:}}=L_{1}\exp{\left(d\left(3+\frac{4\log{\log{T_{0}}}}{\log{T_{0}}}\right)\right)}\max\left\{1,L_{0}^{-\frac{2d}{\log{T_{0}}}}\right\}\times$
×$\times\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}\max\left\{1,K_{1}^{\frac{3d}{\log{T_{0}}}}\right\}\max\left\{L_{0}^{\frac{2}{2-\sigma_{1}}},L_{0}^{\frac{2}{2-\sigma_{2}}}\right\}\times$
×$\times\max\left\{K_{3}^{\frac{2\sigma_{1}-1}{2-\sigma_{1}}},K_{3}^{\frac{2\sigma_{2}-1}{2-\sigma_{2}}}\right\}\max\left\{K_{1}^{\frac{3(1-\sigma_{1})}{2-\sigma_{1}}},K_{1}^{\frac{3(1-\sigma_{2})}{2-\sigma_{2}}}\right\}$
and
$\mathfrak{A}_{2}$ $=\mathfrak{A}_{2}(\sigma_{1},\sigma_{2},T_{0},d,L_{0},L_{2},K_{2},K_{3})\mathrel{\mathop{:}}=L_{2}e^{3d}\max\left\{1,K_{2}^{\frac{3d}{\log{T_{0}}}}\right\}\max\left\{1,L_{0}^{-\frac{2d}{\log{T_{0}}}}\right\}\times$
×$\times\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}\max\left\{L_{0}^{\frac{2}{2-\sigma_{1}}},L_{0}^{\frac{2}{2-\sigma_{2}}}\right\}\max\left\{K_{3}^{\frac{2\sigma_{1}-1}{2-\sigma_{1}}},K_{3}^{\frac{2\sigma_{2}-1}{2-\sigma_{2}}}\right\}\max\left\{K_{2}^{\frac{3(1-\sigma_{1})}{2-\sigma_{1}}},K_{2}^{\frac{3(1-\sigma_{2})}{2-\sigma_{2}}}\right\},$
where $K_{1}=K_{1}(T_{0},A_{0},\kappa,h)$, $K_{2}=K_{2}(T_{0},A_{0},\kappa)$ and $K_{3}=K_{3}(T_{0},\kappa,h,\delta)$ are defined by (37), (38) and (36), respectively, while $L_{0}=L_{0}(T_{0},H_{0},\kappa,h,\delta)$, $L_{1}=L_{1}(T_{0},K_{1},K_{3},h,\delta)$ and $L_{2}=L_{2}(T_{0},K_{2},K_{3},h,\delta)$ are defined by (35), (39) and (40), respectively. Moreover,
$\mathfrak{A}_{3}(\sigma_{1},T_{0},d)\mathrel{\mathop{:}}=\frac{1}{\pi\log{2}}\left(\frac{3-2\sigma_{1}}{2d}+\frac{1}{\log{T_{0}}}\right),\quad\mathfrak{A}_{4}(\sigma_{1},T_{0},d)\mathrel{\mathop{:}}=\frac{3}{2\log{2}}\left(\frac{3-2\sigma_{1}}{2d}+\frac{1}{\log{T_{0}}}\right),$
$\mathfrak{A}_{5}(\sigma_{1},T_{0},H_{0},d,h)\mathrel{\mathop{:}}=\frac{C_{6}}{2\pi}\left(\frac{3-2\sigma_{1}}{2d}+\frac{1}{\log{T_{0}}}\right)+\frac{C_{5}}{2\pi d},$
where $C_{5}=C_{5}(hT_{0},h)$ and $C_{6}=C_{6}(hT_{0},H_{0})$ are defined by (33) and (8), respectively.
Proof of Corollary 1.
We are using Theorem 2 with $T_{0}=H_{0}=H=H_{\mathrm{RH}}$, $h=1$ and $0<d\leq 0.89$. Then
$N(\sigma,T)$ $\leq\frac{1}{2\pi d}\left(\mathfrak{A}_{1}+\mathfrak{A}_{2}\cdot\left(\log{H_{\mathrm{RH}}}\right)^{-\frac{6(1-\sigma_{2})}{2-\sigma_{2}}}\right)T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}$
$+\left(\mathfrak{A}_{3}+\mathfrak{A}_{4}\right)\log^{2}{T}+\left(\mathfrak{A}_{3}\log{H_{\mathrm{RH}}}+\mathfrak{A}_{5}\right)\log{T}$ (10)
under the conditions of Theorem 1. Note that the functions $\mathfrak{A}_{n}$, $n\in\{1,\ldots,5\}$, are decreasing in $\sigma_{1}$. This implies that (10) holds for $\sigma\in(1/2+d/\log{T},0.53125]$ with $\sigma_{1}=1/2$ and $\sigma_{2}=0.53125$. The values for the parameters $d$, $\kappa$, $\delta$ and $A_{0}$ were obtained (see Table 1) in a way that the main term in (10) is the smallest possible. Because the conditions of Theorem 1 are then satisfied, this implies
$N(\sigma,T)\leq 5.36\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+9.461\log^{2}{T}+167.8\log{T}$
for $\sigma\in(1/2+0.28/\log{T},0.53125]$ and $T\geq H_{\mathrm{RH}}$. The above estimate is true also for $\sigma\in[1/2,1/2+0.28/\log{T}]$ since
$N(T)\leq\frac{T}{2\pi}\log{\frac{T}{2\pi e}}+0.43\log{T}\leq 5.36\exp{\left(\frac{-0.56}{\frac{3}{2}-\frac{0.28}{\log{H_{\mathrm{RH}}}}}\right)}T\log^{3}{T}+9.461\log^{2}{T}+167.8\log{T}$ (11)
holds for all $T\geq H_{\mathrm{RH}}$ and
$T\exp{\left(\frac{-0.56}{\frac{3}{2}-\frac{0.28}{\log{T_{0}}}}\right)}\leq T^{\frac{3(1-\sigma)}{2-\sigma}}$
holds for $\sigma\in[1/2,1/2+0.28/\log{T}]$ and $T\geq H_{\mathrm{RH}}$, while the first inequality in (11) follows by [BW24]. This proves (3) for $\sigma\in[1/2,0.53125]$. For the other ranges, the procedure is similar: the values for the parameters $d$, $\kappa$, $\delta$ and $A_{0}$ are obtained for each interval so that the main term in (10) is the smallest possible; see Table 1. In each case, the conditions of Theorem 1 are satisfied, so (10) holds. The proof is thus complete. ∎
2.2. The fourth power moment estimate
The proof of Theorem 2 will be presented in Section 5. The $2k$-th power moment $\mathcal{M}_{k}(A,B)$ is defined by (4).
Theorem 2.
Assume the following notation and conditions:
  1. (1)
    $T\geq T_{0}\geq 55$;
  2. (2)
    $\bm{c}=(c_{1},c_{2})$ with $c_{1}\in(-1,-1/2)$ and $c_{2}\in(0,1/2)$;
  3. (3)
    $\bm{\sigma}=\left(\sigma_{1},\bm{\sigma^{\prime}},\bm{\sigma^{\prime\prime}}\right)$ with $\sigma_{1}\in[0,1/2)$, $\bm{\sigma^{\prime}}=(\sigma_{2},\sigma_{3})$ with $\sigma_{2}\in(1/2,1]$ and $\sigma_{3}\in[0,1/2)$, and $\bm{\sigma^{\prime\prime}}=(\sigma_{4},\sigma_{5})$ with $\sigma_{4}\in(1/2,|c_{1}|)$ and $\sigma_{5}\in(1/2-c_{2},1/2)$;
  4. (4)
    $\bm{\mathfrak{a}}=\left(\mathfrak{a}_{1},\mathfrak{a}_{2}\right)$ with $\mathfrak{a}_{1}>0$ and $\mathfrak{a}_{2}>1$, and $\bm{\mathfrak{b}}=\left(\mathfrak{b}_{1},\mathfrak{b}_{2}\right)$ with $\mathfrak{b}_{1}>0$ and $\mathfrak{b}_{2}>1$;
  5. (5)
    $T-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}\geq T_{0}-\mathfrak{a}_{1}\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}}\geq 2$, $\left(\log{T}\right)^{\mathfrak{a}_{2}}/T\leq\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}}/T_{0}$ and $\mathfrak{a}_{1}\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}-1}\geq 1$;
  6. (6)
    $T-\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}\geq T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\geq 2$, $\left(\log{T}\right)^{\mathfrak{b}_{2}}/T\leq\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}/T_{0}$ and $\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}-1}\geq 2$.
Then
$\left|\mathcal{M}_{2}(T,2T)-\frac{1}{2\pi^{2}}T\log^{4}{T}\right|\leq\mathfrak{F}_{1}\left(T_{0},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{T}$ (12)
and
$\mathcal{M}_{2}(T,2T)\leq\frac{1}{\pi^{2}}T\log^{4}{T}+\left(2\sum_{n=3}^{6}\left(\frac{\mathcal{J}_{n}}{\pi^{2}}\log{T}+2\mathcal{J}_{1}\mathcal{J}_{n}\right)^{\frac{1}{2}}+\mathfrak{F}_{2}\left(T_{0},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\right)T\log^{3}{T},$ (13)
where
$\mathfrak{F}_{1}\left(T_{0},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\mathcal{J}_{0}+\mathcal{J}_{1}+\mathcal{J}_{2,1}+2\mathcal{J}_{2,2}+\sum_{n=3}^{6}\mathcal{J}_{n}+2\mathop{\sum\sum}_{3\leq n<m\leq 6}\left(\mathcal{J}_{n}\mathcal{J}_{m}\right)^{\frac{1}{2}}$
and
$\mathfrak{F}_{2}\left(T_{0},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)=\mathcal{J}_{0}+2\mathcal{J}_{1}+\sum_{n=3}^{6}\mathcal{J}_{n}+2\mathop{\sum\sum}_{3\leq n<m\leq 6}\left(\mathcal{J}_{n}\mathcal{J}_{m}\right)^{\frac{1}{2}}.$
Here, $\mathcal{J}_{n}=\mathcal{J}_{n}(T_{0})$ for $n\in\{1,3,4\}$ are defined by (75), (106) and (107), respectively, while $\mathcal{J}_{0}=\mathcal{J}_{0}\left(T_{0},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$, $\mathcal{J}_{2,1}=\mathcal{J}_{2,1}(T_{0},\sigma_{1})$, $\mathcal{J}_{2,2}=\mathcal{J}_{2,2}\left(T_{0},\bm{\sigma^{\prime}},\bm{\sigma^{\prime\prime}},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$, $\mathcal{J}_{5}=\mathcal{J}_{5}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathcal{J}_{6}=\mathcal{J}_{6}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (116), (78), (105), (108) and (109), respectively.
Corollary 4.
Assume the notation and conditions of Theorem 2. Then
$\mathcal{M}_{2}(T_{0},T)\leq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}+\mathfrak{F}_{1}\left(T_{0},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{\frac{T}{2}}+\mathcal{M}_{2}(T_{0},2T_{0})$ (14)
and
$\mathcal{M}_{2}(T_{0},T)\geq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}-\left(\frac{2\log{2}}{\pi^{2}}+\mathfrak{F}_{1}\left(T_{0},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\right)T\log^{3}{\frac{T}{2}}-\frac{T_{0}}{\pi^{2}}\log^{4}{\frac{T}{2}}.$ (15)
Furthermore,
$\mathcal{M}_{2}(T_{0},T)\leq\frac{1}{\pi^{2}}T\log^{4}{\frac{T}{2}}+\left(2\sum_{n=3}^{6}\left(\frac{\mathcal{J}_{n}}{\pi^{2}}\log{\frac{T}{2}}+2\mathcal{J}_{1}\mathcal{J}_{n}\right)^{\frac{1}{2}}+\mathfrak{F}_{2}\left(T_{0},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\right)T\log^{3}{\frac{T}{2}}+\mathcal{M}_{2}(T_{0},2T_{0}).$ (16)
Proof.
It is clear that all three estimates hold for $T\in[T_{0},2T_{0}]$, so we can assume $T\geq 2T_{0}$. Let $n_{0}\geq 1$ be an integer such that $2^{-n_{0}}T\geq T_{0}$ but $2^{-n_{0}-1}T<T_{0}$. We then have
$\mathcal{M}_{2}(T_{0},T)=\mathcal{M}_{2}\left(T_{0},\frac{T}{2^{n_{0}}}\right)+\sum_{n=0}^{n_{0}-1}\mathcal{M}_{2}\left(\frac{T}{2^{n+1}},\frac{T}{2^{n}}\right)$
by a dyadic partition. We can use the estimates from Theorem 2 to bound the above sum due to our conditions. Then the upper bounds from Corollary 4 immediately follow. The lower bound is derived as follows. Using $(1-x)^{4}-1\geq-4x$, which is valid for $x\geq 0$, we obtain
$\frac{1}{2\pi^{2}}\sum_{n=0}^{n_{0}-1}\frac{T}{2^{n+1}}\left(\log{\frac{T}{2^{n+1}}}\right)^{4}$ $\geq\frac{1}{2\pi^{2}}\left(\sum_{n=0}^{n_{0}-1}\frac{1}{2^{n+1}}\right)T\log^{4}{\frac{T}{2}}-\frac{2\log{2}}{\pi^{2}}\left(\sum_{n=0}^{n_{0}-1}\frac{n}{2^{n+1}}\right)T\log^{3}\frac{T}{2}$
$\geq\frac{1}{2\pi^{2}}\left(1-\frac{2T_{0}}{T}\right)T\log^{4}{\frac{T}{2}}-\frac{2\log{2}}{\pi^{2}}T\log^{3}\frac{T}{2}.$
The result now follows from (12). ∎
Corollary 5.
Let $A_{0}\geq 10^{5}$. Then we have $\mathcal{M}_{2}(A_{0})\leq a_{1}(A_{0})$ and $\mathcal{M}_{2}(A_{0},T)\leq a_{2}(A_{0})T\log^{4}{T}$ for every $T\geq A_{0}$, where
$a_{1}(A_{0})\mathrel{\mathop{:}}=\frac{1}{2\pi^{2}}A_{0}\log^{4}\frac{A_{0}}{2}+48.801A_{0}\log^{3}\frac{A_{0}}{2}+3.0592\cdot 10^{10}$
and
$a_{2}(A_{0})\mathrel{\mathop{:}}=\frac{1}{2\pi^{2}}\left(1-\frac{\log{2}}{\log{A_{0}}}\right)^{4}+\frac{48.801}{\log{A_{0}}}\left(1-\frac{\log{2}}{\log{A_{0}}}\right)^{3}+\frac{2.1817\cdot 10^{10}}{A_{0}\log^{4}{A_{0}}}.$
Proof.
The first estimate follows by the second inequality in (9) of Corollary 2. Because $\mathcal{M}_{2}\left(A_{0},T\right)\leq\mathcal{M}_{2}\left(10^{5},T\right)$, we have $\mathcal{M}_{2}(A_{0},T)\leq a_{2}(T)T\log^{4}{T}$ by (18), see the proof of Corollary 2. Derivative analysis shows that $a_{2}(T)$ is a decreasing function for $T\geq 10^{5}$, implying that $a_{2}(T)\leq a_{2}(A_{0})$. This gives the second estimate. ∎
Proof of Corollary 2.
Assume $T\geq T_{0}^{\prime}=T_{0}\geq 55$, and also that the conditions from Theorem 2 are satisfied. By (16) we have
$\mathcal{M}_{2}(T_{0}^{\prime},T)\leq\frac{1}{\pi^{2}}T\log^{4}{\frac{T}{2}}+\mathcal{C}_{1}\left(T_{0}^{\prime},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{\frac{7}{2}}{\frac{T}{2}}+\mathcal{M}_{2}(T_{0}^{\prime},2T_{0}^{\prime})$
for $T\geq T_{0}^{\prime}$, where
$\mathcal{C}_{1}\left(T_{0}^{\prime},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=2\sum_{n=3}^{6}\left(\frac{\mathcal{J}_{n}}{\pi^{2}}+\frac{2\mathcal{J}_{1}\mathcal{J}_{n}}{\log{(T_{0}^{\prime}/2)}}\right)^{\frac{1}{2}}+\frac{\mathfrak{F}_{2}\left(T_{0}^{\prime},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)}{\log^{1/2}{(T_{0}^{\prime}/2)}}.$
We are taking $T_{0}^{\prime}=3\cdot 10^{3}$ because for this value the calculations via numerical integration are still reasonably fast. We obtain $7.031\cdot 10^{5}<\mathcal{M}_{2}(T_{0}^{\prime})<7.032\cdot 10^{5}$ and $1.249\cdot 10^{6}<\mathcal{M}_{2}(T_{0}^{\prime},2T_{0}^{\prime})<1.25\cdot 10^{6}$. For the choices $\bm{c}=(-0.923364,0.161176)$, $\bm{\mathfrak{a}}=(0.531241,1.52906)$ and $\bm{\mathfrak{b}}=(0.754804,1.68921)$, the value of $\mathcal{C}_{1}(T_{0}^{\prime},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}})$ is $20.7225$, which is as small [^7] According to the method FindMinimum from Mathematica. It should also be noted that all computations in this paper were performed with this software system. as possible. We can also see that the conditions (5) and (6) from Theorem 2 are satisfied by this choice of parameters. Therefore, Corollary 4 implies that
$\mathcal{M}_{2}\left(T_{0}^{\prime},T\right)\leq\frac{1}{\pi^{2}}T\log^{4}{\frac{T}{2}}+20.7225\cdot T\log^{\frac{7}{2}}{\frac{T}{2}}+1.25\cdot 10^{6}$ (17)
for $T\geq T_{0}^{\prime}=3\cdot 10^{3}$. Inequality (8) now easily follows from (17) since $\mathcal{M}_{2}(T)=\mathcal{M}_{2}\left(T_{0}^{\prime},T\right)+\mathcal{M}_{2}\left(T_{0}^{\prime}\right)$.
Assume $T\geq T_{0}^{\prime\prime}=T_{0}\geq T_{0}^{\prime}=3\cdot 10^{3}$, and also that the conditions from Theorem 2 are satisfied. Note that $\mathcal{M}_{2}\left(T_{0}^{\prime\prime},2T_{0}^{\prime\prime}\right)\leq\mathcal{M}_{2}\left(T_{0}^{\prime},2T_{0}^{\prime\prime}\right)$. By (14) and (17) we have
$\mathcal{M}_{2}\left(T_{0}^{\prime\prime},T\right)\leq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}+\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{\frac{T}{2}}+\mathcal{C}_{2}\left(T_{0}^{\prime\prime}\right),$
where
$\mathcal{C}_{2}\left(T_{0}^{\prime\prime}\right)\mathrel{\mathop{:}}=\frac{2}{\pi^{2}}T_{0}^{\prime\prime}\log^{4}{T_{0}^{\prime\prime}}+20.7225\left(2T_{0}^{\prime\prime}\log^{\frac{7}{2}}{T_{0}^{\prime\prime}}\right)+1.25\cdot 10^{6}.$
We are taking $T_{0}^{\prime\prime}=10^{5}$. Finding a global minimum for $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ is a computationally very complex problem, since we need to optimize 11 parameters that satisfy the conditions (2)–(4) from Theorem 2. To simplify the process, we used randomization; we generate $50$ trials where for each trial we randomly choose $100$ sets of values for these parameters, then we select an element from this set such that $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ is minimal, and finally the “minimal” set from all trials. The values we obtained with such a method are $\bm{c}=(-0.938,0.194)$, $\bm{\sigma}=(0.368,(0.718,0.435),(0.629,0.3654))$, $\bm{\mathfrak{a}}=(1.99,1.44)$, and $\bm{\mathfrak{b}}=(1.33,1.45)$, giving $\mathfrak{F}_{1}\left(T_{0}^{\prime\prime},\bm{c},\bm{\sigma},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\leq 48.801$. The conditions (5) and (6) from Theorem 2 are clearly satisfied by this choice of parameters. Trivially, $\mathcal{C}_{2}\left(T_{0}^{\prime\prime}\right)\leq 2.1817\cdot 10^{10}$. Therefore,
$\mathcal{M}_{2}\left(10^{5},T\right)\leq\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}+48.801\cdot T\log^{3}{\frac{T}{2}}+2.1817\cdot 10^{10}$ (18)
for $T\geq 10^{5}$. The second inequality in (9) now easily follows from (18) since $\mathcal{M}_{2}\left(10^{5}\right)\leq 8.775\cdot 10^{9}$ by (8). The first inequality in (9) follows from (15) since $\mathcal{M}_{2}(T)\geq\mathcal{M}_{2}\left(10^{5},T\right)$. The proof of Corollary 2 is thus complete. ∎
Proof of Corollary 3.
We have
$\mathcal{M}_{1}(T)\leq\left(1+\frac{0.548}{\log{T_{0}}}\right)T\log{T},\quad\mathcal{M}_{2}(T)\leq\left(\frac{1}{2\pi^{2}}+\frac{241}{\log{T_{0}}}\right)T\log^{4}{T}$
for $T\geq T_{0}\geq 10^{5}$ by [DZA22, Theorem 3.4] for the first inequality, and by Corollary 2 and derivative analysis for the second inequality. The result now follows by the CBS inequality since then $\mathcal{M}_{2/3}(T)\leq\left(\mathcal{M}_{1}(T)\mathcal{M}_{2}(T)\right)^{1/2}$. ∎
3. Gabriel’s convexity theorem and applications
Gabriel [Gab27, Theorem 2] found an interesting two-variable convexity estimate for integrals of holomorphic functions along vertical lines in the strip, see Lemma 3.
Lemma 3.
Let $f(z)$ be a holomorphic function in the strip $\alpha<\Re\{z\}<\beta$ for $\alpha<\beta$ with continuous extension to the boundary $\Re\{z\}=\alpha$, $\Re\{z\}=\beta$. Assume that
$\lim_{\begin{subarray}{c}|y|\to\infty\\ y\in\mathbb{R}\end{subarray}}|f(x+\mathrm{i}y)|=0$
uniformly in $x\in[\alpha,\beta]$. For $c>0$ define
$G_{f}(x,c)\mathrel{\mathop{:}}=\left(\int_{-\infty}^{\infty}\left|f(x+\mathrm{i}y)\right|^{\frac{1}{c}}\mathrm{d}y\right)^{c}.$ (19)
Then
$G_{f}(x,w_{1}\lambda+w_{2}\mu)\leq G_{f}^{w_{1}}(\alpha,\lambda)G_{f}^{w_{2}}(\beta,\mu)$ (20)
for $x\in[\alpha,\beta]$, where
$w_{1}=w_{1}(x,\alpha,\beta)\mathrel{\mathop{:}}=\frac{\beta-x}{\beta-\alpha},\quad w_{2}=w_{2}(x,\alpha,\beta)\mathrel{\mathop{:}}=\frac{x-\alpha}{\beta-\alpha},$ (21)
and $\lambda$ and $\mu$ are arbitrary positive numbers.
Proof.
See, e.g., [KV92, pp. 381–384]. ∎
This result should be compared with [HIP27, Theorem 7], where (20) with $\lambda=\mu=1/p$ was proved under different growth conditions on $f(z)$.
Let $y\mapsto f(x+\mathrm{i}y)$ be a continuous function for some fixed $x\in\mathbb{R}$ and $A\leq y\leq B$. Then
$I_{f}(x,A,B,c)\mathrel{\mathop{:}}=\int_{A}^{B}\left|f(x+\mathrm{i}y)\right|^{\frac{1}{c}}\mathrm{d}y$ (22)
exists for every $c>0$. The next theorem shows how we can replace, under certain mild conditions on $f(z)$, the infinite integral (19) with the finite integral (22), and still more or less preserve the form of Gabriel’s inequality (20).
Theorem 4.
Let $1/2\leq\alpha<\beta$ and consider a holomorphic function $f(z)$ in $\{z\in\mathbb{C}\colon\alpha<\Re\{z\}<\beta\}\setminus\{1\}$ with continuous extension to the vertical lines $\Re\{z\}=\alpha$ and $\Re\{z\}=\beta$ such that:
  1. (1)
    If $\beta>1$, then $f(z)$ has a pole of order $m\geq 0$ at $z=1$. If $\beta\leq 1$, we set $m=0$.
  2. (2)
    For any fixed $d>0$, $\lim_{|y|\to\infty,y\in\mathbb{R}}e^{-dy^{2}}|f(x+\mathrm{i}y)|=0$ uniformly in $x\in[\alpha,\beta]$.
  3. (3)
    For $x\in\{\alpha,\beta\}$ and for any fixed $a\geq 0$, $b>0$ and $c>0$, $\lim_{|u|\to\infty,u\in\mathbb{R}}e^{-u^{2}}I_{f}(x,a,bu,c)=0$.
Also, let $T>H\geq 0$, $\kappa>0$, $A_{1}\geq 0$, and $A_{2}\geq 0$, and let $\lambda$ and $\mu$ be arbitrary positive numbers. Then
$\left(I_{f}\left(x,-T,-H,w_{1}\lambda+w_{2}\mu\right)+I_{f}\left(x,H,T,w_{1}\lambda+w_{2}\mu\right)\right)^{w_{1}\lambda+w_{2}\mu}\leq\left(\frac{(x-1)^{2}+H^{2}}{x^{2}+H^{2}}\right)^{-\frac{m}{2}}\times\\ \times\exp{\left(\kappa\left(1+\left(\frac{\beta-\alpha}{2T}\right)^{2}\right)\right)}\left(\mathcal{J}(\alpha,A_{1},\lambda)\right)^{w_{1}\lambda}\left(\mathcal{J}(\beta,A_{2},\mu)\right)^{w_{2}\mu}$ (23)
for $x\in[\alpha,\beta]$, where
$\mathcal{J}(a,b,c)\mathrel{\mathop{:}}=I_{f}(a,-b,b,c)+\int_{\frac{\kappa}{c}\left(\frac{b}{T}\right)^{2}}^{\infty}e^{-u}\left(I_{f}\left(a,-T\sqrt{\frac{cu}{\kappa}},-b,c\right)+I_{f}\left(a,b,T\sqrt{\frac{cu}{\kappa}},c\right)\right)\mathrm{d}u,$ (24)
$w_{1}=w_{1}(x,\alpha,\beta)$ and $w_{2}=w_{2}(x,\alpha,\beta)$ are given by (21), and $H>0$ if $x=1$ and $m\neq 0$.
Proof.
Write $z=x+\mathrm{i}y$ for $(x,y)\in\mathbb{R}^{2}$ and define
$g(z)\mathrel{\mathop{:}}=\left(\frac{z-1}{z}\right)^{m}\exp{\left(\kappa\left(\frac{z}{T}\right)^{2}\right)},\quad\phi(z)\mathrel{\mathop{:}}=g(z)f(z),$
where $m$ is from the condition (1) of Theorem 4. It is clear that $\phi(z)$ is a holomorphic function in the strip $\alpha<x<\beta$ with continuous extension to the vertical lines $x=\alpha$ and $y=\beta$. By a slight modification of [KLN18, Lemma 3.7] we have $|g(z)|\leq\omega_{1}(x)e^{-\kappa(|y|/T)^{2}}$ for $x\geq 1/2$, $y\in\mathbb{R}$, and $|g(z)|\geq\omega_{2}(x)$ for $x\geq 1/2$, $|y|\in[H,T]$, where
$\omega_{1}(x)\mathrel{\mathop{:}}=\exp{\left(\kappa\left(\frac{x}{T}\right)^{2}\right)},\quad\omega_{2}(x)\mathrel{\mathop{:}}=\left(\frac{(x-1)^{2}+H^{2}}{x^{2}+H^{2}}\right)^{\frac{m}{2}}\exp{\left(\kappa\left(\frac{x}{T}\right)^{2}-\kappa\right)}.$
This also shows that $\lim_{|y|\to\infty}\left|\phi(x+\mathrm{i}y)\right|=0$, uniformly in $x\in[\alpha,\beta]$, due to the condition (2). Therefore, Lemma 3 can be applied for $\phi(z)$. We are going to provide upper and lower bounds for $G_{\phi}(x,c)$, where this function is defined by (19). For $A\geq 0$, $x\in\{\alpha,\beta\}$ and $c>0$ we have
$G_{\phi}(x,c)$ $\leq\omega_{1}(x)\left(\int_{-\infty}^{\infty}\exp{\left(-\frac{\kappa}{c}\left(\frac{y}{T}\right)^{2}\right)}\left|f(x+\mathrm{i}y)\right|^{\frac{1}{c}}\mathrm{d}y\right)^{c}$
$=\omega_{1}(x)\left(\int_{-\infty}^{\infty}e^{-v^{2}}\left(\frac{\mathrm{d}}{\mathrm{d}v}I_{f}\left(x,A,T\sqrt{\frac{c}{\kappa}}v,c\right)\right)\mathrm{d}v\right)^{c}$ (25)
$=\omega_{1}(x)\left(\int_{-\infty}^{\infty}2ve^{-v^{2}}I_{f}\left(x,A,T\sqrt{\frac{c}{\kappa}}v,c\right)\mathrm{d}v\right)^{c}$ (26)
$\leq\omega_{1}(x)\left(\mathcal{J}(x,A,c)\right)^{c},$ (27)
where $\mathcal{J}(a,b,c)$ is defined by (24). Equality (25) follows by taking $y=Tv\sqrt{c/\kappa}$ and noticing that
$\frac{\mathrm{d}}{\mathrm{d}v}I_{f}\left(x,A,T\sqrt{\frac{c}{\kappa}}v,c\right)=T\sqrt{\frac{c}{\kappa}}\left.\frac{\partial}{\partial{B}}I_{f}(x,A,B,c)\right|_{B=T\sqrt{\frac{c}{\kappa}}v}=T\sqrt{\frac{c}{\kappa}}\left|f\left(x+\mathrm{i}T\sqrt{\frac{c}{\kappa}}v\right)\right|^{\frac{1}{c}},$
while (26) follows by integration by parts and the condition (3). Inequality (27) follows by splitting the integral, taking $v^{2}=u$ and observing that
$I_{f}\left(x,-T\sqrt{\frac{cu}{\kappa}},A,c\right)=I_{f}\left(x,-T\sqrt{\frac{cu}{\kappa}},-A,c\right)+I_{f}\left(x,-A,A,c\right),$
and also that $I_{f}(x,A,T\sqrt{cu/\kappa},c)\leq 0$ and $I_{f}(x,-T\sqrt{cu/\kappa},-A,c)\leq 0$ for $u\in[0,(\kappa/c)(A/T)^{2}]$. On the other hand,
$G_{\phi}(x,c_{1})\geq\omega_{2}(x)\left(\left(\int_{-T}^{-H}+\int_{H}^{T}\right)\left|f(x+\mathrm{i}y)\right|^{\frac{1}{c_{1}}}\mathrm{d}y\right)^{c_{1}}=\omega_{2}(x)\left(I_{f}(x,-T,-H,c_{1})+I_{f}(x,H,T,c_{1})\right)^{c_{1}}$ (28)
for $x\in[\alpha,\beta]$, $c_{1}>0$ and $T>H>0$ if $x=1$ and $m\neq 0$. To obtain an upper bounds for $G_{\phi}(\alpha,\lambda)$ and $G_{\phi}(\beta,\mu)$ from (20), we are using (27) with $x=\alpha$, $c=\lambda$ and $A=A_{1}$ for the former function, and with $x=\beta$, $c=\mu$ and $A=A_{2}$ for the latter function. Lower bound for $G_{\phi}(x,w_{1}\lambda+w_{2}\mu)$ from (20) is given by (28) for $c_{1}=w_{1}\lambda+w_{2}\mu$. The claim of Theorem 4 now follows by Lemma 3 and
$\alpha^{2}(\beta-x)+\beta^{2}(x-\alpha)-x^{2}(\beta-\alpha)\leq\frac{1}{4}(\beta-\alpha)^{3}.$
The latter is true since the function on the left-hand side of the above inequality in the variable $x\in[\alpha,\beta]$ attains the maximum value at $x=(\alpha+\beta)/2$. ∎
Our main application of Theorem 4 is to the mollifier $f_{X}(s)$ of the Riemann zeta-function. Let $X\geq 1$ be some parameter, independent of a complex variable $s=\sigma+\mathrm{i}t$, and define
$f_{X}(s)\mathrel{\mathop{:}}=\zeta(s)M_{X}(s)-1,\quad M_{X}(s)\mathrel{\mathop{:}}=\sum_{n\leq X}\frac{\mu(n)}{n^{s}},$ (29)
where $\mu(n)$ is the Möbius function. Clearly, $f_{X}(s)$ is a holomorphic function on $\mathbb{C}\setminus\{1\}$ with a simple pole at $s=1$, $f_{X}(\bar{s})=\overline{f_{X}(s)}$, and also that $f_{X}(s)\ll X|t|$ for $\sigma\geq 0$ and large $|t|$. For $0\leq A\leq B$ and $c>0$ define
$F_{X}(\sigma,A,B,c)\mathrel{\mathop{:}}=\int_{A}^{B}\left|f_{X}(\sigma+\mathrm{i}t)\right|^{\frac{1}{c}}\mathrm{d}t.$
This function is well-defined for all $\sigma\in\mathbb{R}$ if $A>0$, and for all $\sigma\in\mathbb{R}\setminus\{1\}$ if $A=0$. The next corollary is essentially a slightly simplified inequality (23) for the mollifier $f_{X}(s)$. This result is crucial in the proof of Theorem 1.
Corollary 6.
Let 1/$21/2\leq\alpha<1<\beta\leq 3/$ 2, $T\geq H>1$, $\kappa>0$, $A\geq 0$, and also let $\lambda$ and $\mu$ be arbitrary positive numbers. Then
$\left(F_{X}\left(\sigma,H,T,w_{1}\lambda+w_{2}\mu\right)\right)^{w_{1}\lambda+w_{2}\mu}\leq\left(1-\frac{1}{H}\right)^{-1}\exp{\left(\kappa\left(1+\left(\frac{\beta-\alpha}{2T}\right)^{2}\right)\right)}\times\\ \left(F_{X}(\alpha,0,A,\lambda)+\int_{\frac{\kappa}{\lambda}\left(\frac{A}{T}\right)^{2}}^{\infty}e^{-u}F_{X}\left(\alpha,A,T\sqrt{\frac{\lambda u}{\kappa}},\lambda\right)\mathrm{d}u\right)^{w_{1}\lambda}\left(\int_{0}^{\infty}e^{-u}F_{X}\left(\beta,0,T\sqrt{\frac{\mu u}{\kappa}},\mu\right)\mathrm{d}u\right)^{w_{2}\mu}$ (30)
for $\sigma\in[\alpha,\beta]$, where $w_{1}=w_{1}(\sigma,\alpha,\beta)$ and $w_{2}=w_{2}(\sigma,\alpha,\beta)$ are given by (21).
Proof.
Inequality (30) is obviously true in the case $H=T$, so we can assume that $T>H>1$. The claim then follows by Theorem 4 for $f=f_{X}(s)$, $m=1$, $A_{1}=A$ and $A_{2}=0$ since $f_{X}(\bar{s})=\overline{f_{X}(s)}$,
$\frac{(\sigma-1)^{2}+H^{2}}{\sigma^{2}+H^{2}}\geq 1-\frac{8}{9+(2H)^{2}}\geq\left(1-\frac{1}{H}\right)^{2}$
for $\sigma\in[1/2,3/2]$ and $H>1$. Note that all the conditions of Theorem 4 are satisfied by the properties of $f_{X}(s)$ mentioned before. ∎
Theorem 4 can also be used to prove a convexity property for the power moments of holomorphic functions in a strip.
Corollary 7.
Let $T\geq T_{0}\geq 2$ and $11/2\leq\alpha<\beta<$ 1. Assume that $f(z)$ is a holomorphic function in the strip $x=\Re\{z\}\in(\alpha,\beta)$ with continuous extension to the boundary, such that the conditions (2) and (3) from Theorem 4 are satisfied. Assume also that
$\int_{-T}^{T}\left|f(\alpha+\mathrm{i}y)\right|^{b_{1}}\mathrm{d}y\leq M_{1}T^{a_{1}}\left(\log{T}\right)^{c_{1}},\quad\int_{-T}^{T}\left|f(\beta+\mathrm{i}y)\right|^{b_{2}}\mathrm{d}y\leq M_{2}T^{a_{2}}\left(\log{T}\right)^{c_{2}}$
for all $T\geq T_{0}$, where $M_{i}>0$, $a_{i}>0$, $b_{i}\geq 1$ and $c_{i}\geq 0$, $i\in\{1,2\}$, are some constants that may depend on $f(z)$. Then
$\int_{-T}^{T}\left|f(x+\mathrm{i}y)\right|^{\mu(b_{1},b_{2};x)}\mathrm{d}y\leq M(x)T^{\mu(a_{1},a_{2};x)}\left(\log{T}\right)^{\mu(c_{1},c_{2};x)}$
for all $T\geq T_{0}$ and $x\in[\alpha,\beta]$, where
$\mu(k,l;x)\mathrel{\mathop{:}}=\frac{kb_{2}(\beta-x)+lb_{1}(x-\alpha)}{b_{2}(\beta-x)+b_{1}(x-\alpha)}$
and, for $\kappa>0$,
$M(x)\mathrel{\mathop{:}}=\exp{\left(\kappa\mu(b_{1},b_{2};x)\left(1+\left(\frac{\beta-\alpha}{2T_{0}}\right)^{2}\right)\right)}C_{1}^{\mu(1,0;x)}C_{2}^{\mu(0,1;x)},$
$C_{i}\mathrel{\mathop{:}}=M_{i}\left(1+\int_{0}^{\infty}e^{-u}\left(\frac{u}{\kappa b_{i}}\right)^{\frac{a_{i}}{2}}\left(1+\frac{\log{\left(1+\frac{u}{\kappa b_{i}}\right)}}{2\log{T_{0}}}\right)^{c_{i}}\mathrm{d}u\right).$
Proof.
We are using Theorem 4 with $m=0$, $H=0$, $A_{1}=A_{2}=T_{0}$, $\lambda=1/b_{1}$ and $\mu=1/b_{2}$. By the assumptions we have $I_{f}\left(\alpha,-T_{0},T_{0},1/b_{1}\right)\leq M_{1}T^{a_{1}}\left(\log{T}\right)^{c_{1}}$ and
$I_{f}\left(\alpha,-T\sqrt{\frac{u}{\kappa b_{1}}},-T_{0},\frac{1}{b_{1}}\right)$ $+I_{f}\left(\alpha,T_{0},T\sqrt{\frac{u}{\kappa b_{1}}},\frac{1}{b_{1}}\right)\leq\int_{-T\sqrt{\frac{u}{\kappa b_{1}}}}^{T\sqrt{\frac{u}{\kappa b_{1}}}}\left|f(\alpha+\mathrm{i}y)\right|^{b_{i}}\mathrm{d}y$
$\leq M_{1}T^{a_{1}}\left(\log{T}\right)^{c_{1}}\left(\frac{u}{\kappa b_{1}}\right)^{\frac{a_{1}}{2}}\left(1+\frac{\log\left(1+\frac{u}{\kappa b_{1}}\right)}{2\log{T_{0}}}\right)^{c_{1}}$ (31)
since $u\geq\kappa b_{1}\left(T_{0}/T\right)^{2}$. Both inequalities hold also for $\beta$, $M_{2}$, $a_{2}$, $b_{2}$ and $c_{2}$ in place of $\alpha$, $M_{1}$, $a_{1}$, $b_{1}$ and $c_{1}$, respectively. Observe that the right-hand side of (3) is non-negative for all $u\geq 0$. The result now follows since $1/\left(w_{1}\lambda+w_{2}\mu\right)=\mu(b_{1},b_{2};x)$, $w_{1}\lambda/\left(w_{1}\lambda+w_{2}\mu\right)=\mu(1,0;x)$, $w_{2}\mu/\left(w_{1}\lambda+w_{2}\mu\right)=\mu(0,1;x)$, $k\mu(1,0;x)+l\mu(0,1;x)=\mu(k,l;x)$, and $\mu(1,1;x)=1$. ∎
It should be noted that Corollary 7 also holds for $\beta\geq 1$ if $f(z)$ is holomorphic in the new extended strip. It can be applied to general $L$-functions that belong to the Selberg class or are defined as in [IK04, Chapter 5] since for these functions the growth conditions from Theorem 4 are satisfied by the convexity estimates. In the case of the Riemann zeta-function, Corollary 7 is related to some results from [Ivi03, Sections 8.2 and 8.5] and [Tit86, Section 7.8], and could have some effective applications to the generalized divisor problem [BY24].
4. Proof of Theorem 1
As is customary in the classical theory of zero density estimates, let
$h_{X}(s)\mathrel{\mathop{:}}=1-f_{X}^{2}(s)=\left(2-\zeta(s)M_{X}(s)\right)\zeta(s)M_{X}(s),$
where $f_{X}(s)$ and $M_{X}(s)$ are defined by (29). Note that the zeros of $\zeta(s)$ are also the zeros of $h_{X}(s)$. Let $1/2<\sigma^{\prime}<1<\sigma^{\prime\prime}$ and $0<H\leq T$. Assume that $h_{X}(s)$ does not vanish on the boundary of the open rectangle $\left\{z\in\mathbb{C}\colon\sigma^{\prime}<\Re\{z\}<\sigma^{\prime\prime},H<\Im\{z\}<T\right\}$. Then, by Littlewood’s lemma [Tit86, Section 9.9],
$N(\sigma,T)-N(\sigma,H)\leq\frac{1}{2\pi\left(\sigma-\sigma^{\prime}\right)}\Biggl{(}\int_{H}^{T}\log{\left|h_{X}\left(\sigma^{\prime}+\mathrm{i}t\right)\right|}\mathrm{d}t-\int_{H}^{T}\log{\left|h_{X}\left(\sigma^{\prime\prime}+\mathrm{i}t\right)\right|}\mathrm{d}t\\ +\int_{\sigma^{\prime}}^{\sigma^{\prime\prime}}\arg{h_{X}\left(u+\mathrm{i}T\right)}\mathrm{d}u-\int_{\sigma^{\prime}}^{\sigma^{\prime\prime}}\arg{h_{X}\left(u+\mathrm{i}H\right)}\mathrm{d}u\Biggr{)}$ (32)
for $\sigma\in\left[\sigma^{\prime},\sigma^{\prime\prime}\right]$. Before proceeding to the proof of Theorem 1, we need a few technical results. Corollary 6 is used for the estimation of the first integral, see Lemma 9, where two additional results on the second power moments of $M_{X}(1/2+\mathrm{i}t)$ and $f_{X}(1+\delta/\log{X}+\mathrm{i}t)$ are needed, see Lemmas 5 and 6. Estimates for the rest of the integrals in (32) are provided by Lemmas 7 and 8.
Lemma 5.
Let $T>0$, $X\geq X_{0}\geq 10^{9}$. Then
$\int_{0}^{T}\left|M_{X}\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t\leq\left(C_{1}(X_{0})T+C_{2}(X_{0})X\right)\log{X},$
where
$C_{1}(X_{0})\mathrel{\mathop{:}}=\frac{6}{\pi^{2}}+\frac{b_{2}}{\log X_{0}},\quad C_{2}(X_{0})\mathrel{\mathop{:}}=\frac{\pi m_{0}b_{1}}{\log X_{0}}+\frac{6m_{0}}{\pi X_{0}}+\frac{\pi m_{0}b_{2}}{X_{0}\log X_{0}},$
with $b_{1}\mathrel{\mathop{:}}=0.62$, $b_{2}\mathrel{\mathop{:}}=1.048$, and $m_{0}\mathrel{\mathop{:}}=\sqrt{1+(2/3)\sqrt{6/5}}<1.31541$.
Proof.
See [KLN18, Lemma 4.1]. ∎
Lemma 6.
Let $T>0$, $X\geq X_{0}\geq 10^{9}$, and $\delta>0$. Then
$F_{X}\left(1+\frac{\delta}{\log X},0,T,\frac{1}{2}\right)\leq\left(C_{3}(X_{0},\delta)+\frac{C_{4}(X_{0},\delta)\left(T+\pi m_{0}\right)}{X}\right)\log^{2}{X},$
where
$C_{3}(X_{0},\delta)\mathrel{\mathop{:}}=\frac{\pi m_{0}b_{4}}{2\delta}\left(1+\frac{2\delta}{\log X_{0}}\right)^{2}\exp\left(\frac{2\delta\gamma}{\log X_{0}}\right)$
$C_{4}(X_{0},\delta)\mathrel{\mathop{:}}=\frac{b_{4}}{5\delta e^{\delta}}\left(1+\frac{\delta}{\log X_{0}}\right)^{2}+\frac{b_{3}e^{-2\delta}}{\log^{2}X_{0}},$
with $b_{3}\mathrel{\mathop{:}}=0.605$, $b_{4}\mathrel{\mathop{:}}=0.529$, $m_{0}$ as in Lemma 5, and $\gamma$ is the Euler–Mascheroni constant.
Proof.
See [KLN18, Lemma 4.3]. ∎
Lemma 7.
Let $T\geq H\geq 2\pi m_{0}$, $T\geq T_{0}>0$ and $X=hT$ for $h\geq 10^{9}/T_{0}$. Then
$-\int_{H}^{T}\log{\left|h_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|}\mathrm{d}t\leq C_{5}(hT_{0},h),$
where
$C_{5}(u,h)\mathrel{\mathop{:}}=-\log{\left(1-9\left(1+\frac{2+\gamma+\frac{7}{36u}}{\log{u}}\right)^{2}\frac{\log^{2}{u}}{u}\right)}\frac{\left(\frac{3}{h}+8\pi m_{0}\right)\left(1+\frac{3.31}{\log{u}}\right)}{9\left(1+\frac{2+\gamma+\frac{7}{36u}}{\log{u}}\right)^{2}}\log{u}$ (33)
with $m_{0}$ as in Lemma 5 and $\gamma$ is the Euler–Mascheroni constant.
Proof.
We closely follow the proofs of Corollary 4.11 and Lemma 4.13 from [KLN18]. Note that Corollary 4.11 is valid also for $\eta=1/2$ by the continuation principle. Then $\left|f_{X}(3/2+\mathrm{i}t)\right|\leq b_{6}(X_{0},1/2)$ for $X\geq X_{0}\geq 10^{9}$, where the function $b_{6}(X,\eta)$ is from [KLN18, Equation (4.45)]. Our conditions imply $X_{0}\geq 10^{9}$ by setting $X_{0}=hT_{0}$. Then
$-\log{\left|h_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|}\leq-\frac{\log{\left(1-b_{6}(hT_{0},1/2)^{2}\right)}}{b_{6}(hT_{0},1/2)^{2}}\left|f_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|^{2}.$
Moreover,
$\int_{H}^{T}\left|f_{X}\left(\frac{3}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t\leq\left(\frac{3}{h}+8\pi m_{0}\right)\left(1+\frac{3.31}{\log{(hT_{0})}}\right)\frac{\log^{3}(hT_{0})}{hT_{0}}$
by the proof of [KLN18, Lemma 4.13], where we notice that $T-H+2\pi m_{0}\leq T$ and $b_{11}(X,3)\leq b_{11}(X,2)\leq 1+3.31/\log{X}$ for $X\geq 10^{9}$, where $b_{11}(X,\tau)$ is from [KLN18, Equation (4.68)]. The statement of Lemma 7 now easily follows. ∎
Remark 1.
Note that, for fixed $h$, $0<C_{5}(u,h)\sim\left(3/h+8\pi m_{0}\right)u^{-1}\log^{3}{u}\to 0$ as $u\to\infty$.
Lemma 8.
Let $T\geq H\geq H_{0}\geq 1002$, $T\geq T_{0}\geq e^{2d}$ for $d>0$, $1/2+d/\log{T}\leq\sigma\leq 1$ and $X=hT$ for $h\geq 10^{9}/T_{0}$. Then
$\left|\int_{\sigma-\frac{d}{\log{T}}}^{\frac{3}{2}}\arg h_{X}(u+\mathrm{i}T)\mathrm{d}u-\int_{\sigma-\frac{d}{\log{T}}}^{\frac{3}{2}}\arg h_{X}(u+\mathrm{i}H)\mathrm{d}u\right|\leq\left(\frac{3}{2}-\sigma+\frac{d}{\log{T_{0}}}\right)\times\\ \times\left(C_{6}\left(hT_{0},H_{0}\right)+\frac{2}{\log{2}}\log{(HT)}+\frac{3\pi}{\log{2}}\log{(hT)}\right),$
where
$C_{6}(u,H_{0})$ $\mathrel{\mathop{:}}=220+\frac{\pi}{\log{2}}\log{\left(1+\frac{1}{u^{3/2}}\right)}+\frac{4}{\log{2}}\log{\left(\frac{1}{2\pi}\sqrt{\left(\frac{5}{2H_{0}}\right)^{2}+\left(1+\frac{2}{H_{0}}\right)^{2}}\right)}$
$-\frac{2\pi}{\log{2}}\log{\left(1-9\left(1+\frac{2+\gamma+\frac{7}{36u}}{\log{u}}\right)^{2}\frac{\log^{2}{u}}{u}\right)}$ (34)
and $\gamma$ is the Euler–Mascheroni constant.
Proof.
This follows by the proof of Lemma 4.12 from [KLN18], by using similar arguments as in the proof of Lemma 7, and by several numerical simplifications. ∎
Lemma 9.
Let $T\geq H\geq H_{0}>1$, $T\geq T_{0}\geq e^{e}$, and $\kappa>0$. Also, let $X=hT$ for $h\geq 10^{9}/T_{0}$, and let $w_{1}=w_{1}\left(\sigma,1/2,1+\delta/\log{X}\right)$ and $w_{2}=w_{2}\left(\sigma,1/2,1+\delta/\log{X}\right)$ for $0<\delta\leq\frac{1}{2}\log{(hT_{0})}$ and with $w_{1}$ and $w_{2}$ as in (21). Then
$F_{X}\left(\sigma,H,T,\frac{3}{4}w_{1}+\frac{1}{2}w_{2}\right)$ $\leq L_{0}^{\frac{2}{2-\sigma}}K_{3}^{\frac{2\sigma-1}{2-\sigma}}\left(L_{1}\cdot\left(K_{1}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}}\right.$
$\left.+L_{2}\cdot\left(K_{2}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}\right)$
for $1/2\leq\sigma\leq 1$. Functions $L_{0}$ and $K_{3}$ are defined by
$L_{0}=L_{0}(T_{0},H_{0},\kappa,h,\delta)\mathrel{\mathop{:}}=\frac{\exp{\left(\kappa\left(1+\left(\frac{1}{4T_{0}}\left(1+\frac{2\delta}{\log{(hT_{0})}}\right)\right)^{2}\right)\right)}}{1-H_{0}^{-1}},$ (35)
and
$K_{3}=K_{3}(T_{0},\kappa,h,\delta)\mathrel{\mathop{:}}=\int_{0}^{\infty}e^{-u}\left(C_{3}(hT_{0},\delta)+\frac{C_{4}(hT_{0},\delta)}{h}\left(\sqrt{\frac{u}{2\kappa}}+\frac{\pi m_{0}}{T_{0}}\right)\right)\mathrm{d}u,$ (36)
where $C_{3}(X_{0},\delta)$, $C_{4}(X_{0},\delta)$ and $m_{0}$ are from Lemma 6. Moreover,
$K_{1}=K_{1}(T_{0},A_{0},\kappa,h)\mathrel{\mathop{:}}=\left(2a_{2}(A_{0})\sqrt{\frac{3}{4\kappa}}\right)^{\frac{1}{3}}\int_{0}^{\infty}e^{-u}u^{\frac{1}{6}}\left(1+\frac{\log{\left(1+\frac{3u}{4\kappa}\right)}}{2\log{T_{0}}}\right)^{\frac{4}{3}}\times\\ \times\left(C_{1}(hT_{0})\sqrt{\frac{3u}{4\kappa}}+hC_{2}(hT_{0})\right)^{\frac{2}{3}}\mathrm{d}u+\frac{\left(2a_{1}(A_{0})\right)^{\frac{1}{3}}\left(hC_{2}(hT_{0})+\frac{A_{0}C_{1}(hT_{0})}{T_{0}}\right)^{\frac{2}{3}}}{T_{0}^{\frac{1}{3}}\log^{\frac{4}{3}}{T_{0}}}$ (37)
with $C_{1}(X_{0})$ and $C_{2}(X_{0})$ as in Lemma 5, and $a_{1}(A_{0})$, $a_{2}(A_{0})$ and $A_{0}$ as in Corollary 5, while
$K_{2}=K_{2}(T_{0},A_{0},\kappa)\mathrel{\mathop{:}}=2^{\frac{1}{3}}\sqrt{\frac{3}{4\kappa}}\int_{0}^{\infty}e^{-u}\sqrt{u}\mathrm{d}u+2^{\frac{1}{3}}\frac{4\kappa}{3}\left(\frac{A_{0}}{T_{0}}\right)^{3}.$ (38)
Finally,
$L_{1}=L_{1}(T_{0},K_{1},K_{3},h,\delta)\mathrel{\mathop{:}}=\max\left\{1,K_{1}^{\frac{3\delta}{\log{(hT_{0})}}}\right\}\max\left\{1,K_{3}^{-\frac{3\delta}{\log{(hT_{0})}}}\right\}\times\\ \times\exp{\left(3\delta\max\left\{1,\frac{\log{T_{0}}}{\log{(hT_{0})}}\right\}+4\delta\max\left\{\frac{\log{\log{T_{0}}}}{\log{(hT_{0})}},\frac{\log{\log{T_{0}}}}{\log{T_{0}}}\right\}\right)}$ (39)
and
$L_{2}=L_{2}(T_{0},K_{2},K_{3},h,\delta)\mathrel{\mathop{:}}=\max\left\{1,K_{2}^{\frac{3\delta}{\log{(hT_{0})}}}\right\}\max\left\{1,K_{3}^{-\frac{3\delta}{\log{(hT_{0})}}}\right\}\exp{\left(3\delta\max\left\{1,\frac{\log{T_{0}}}{\log{(hT_{0})}}\right\}\right)},$ (40)
where $K_{1}$, $K_{2}$ and $K_{3}$ are given by (37), (38) and (36), respectively.
Proof.
We are using Corollary 6 with $A=A_{0}$, $\alpha=1/2$, $\beta=1+\delta/\log{X}$, $\lambda=3/4$ and $\mu=1/2$. Firstly, we estimate the first term in the second line of (30). Note that
$F_{X}\left(\alpha,U,V,\lambda\right)\leq 2^{\frac{1}{3}}\mathcal{M}_{2}(U,V)^{\frac{1}{3}}\left(\int_{0}^{V}\left|M_{X}\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t\right)^{\frac{2}{3}}+2^{\frac{1}{3}}(V-U)$ (41)
for $V\geq U\geq 0$ by Minkowski’s and Hölder’s inequalities. Take $U=A_{0}$ and $V=T\sqrt{\lambda u/\kappa}$. Note that $u\geq(\kappa/\lambda)(A/T)^{2}$ due to the domain of integration in (30), implying $V\geq U$. Then
$\mathcal{M}_{2}(U,V)\leq a_{2}(A_{0})\sqrt{\frac{3u}{4\kappa}}\left(1+\frac{\log{\left(1+\frac{3u}{4\kappa}\right)}}{2\log{T_{0}}}\right)^{4}T\log^{4}{T}$
by Corollary 5, and
$\int_{0}^{V}\left|M_{X}\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t\leq\left(C_{1}(hT_{0})\sqrt{\frac{3u}{4\kappa}}+hC_{2}(hT_{0})\right)T\log{(hT)}$
by Lemma 5, where we set $X_{0}=hT_{0}$. Also,
$F_{X}\left(\alpha,0,A,\lambda\right)\leq\left(2a_{1}(A_{0})\right)^{\frac{1}{3}}\left(hC_{2}(hT_{0})+\frac{A_{0}C_{1}(hT_{0})}{T_{0}}\right)^{\frac{2}{3}}T^{\frac{2}{3}}\log^{\frac{2}{3}}{(hT)}+2^{\frac{1}{3}}A_{0}$
by (41) for $U=0$ and $V=A_{0}$, and then again by Corollary 5 and Lemma 5. Therefore,
$F_{X}\left(\alpha,0,A,\lambda\right)+\int_{\frac{\kappa}{\lambda}\left(\frac{A}{T}\right)^{2}}^{\infty}e^{-u}F_{X}\left(\alpha,A,T\sqrt{\frac{\lambda u}{\kappa}},\lambda\right)\mathrm{d}u\leq K_{1}T\log^{\frac{4}{3}}{T}\cdot\log^{\frac{2}{3}}{(hT)}+K_{2}T,$
where $K_{1}$ and $K_{2}$ are defined by (37) and (38), respectively. Secondly, for the second term in the second line of (23) we simply have
$\int_{\frac{\kappa}{\mu}\left(\frac{A}{T}\right)^{2}}^{\infty}e^{-u}F_{X}\left(\beta,0,T\sqrt{\frac{\mu u}{\kappa}},\mu\right)\mathrm{d}u\leq K_{3}\log^{2}(hT)$
by Lemma 6, where $K_{3}$ is defined by (36). Remembering that $(a+b)^{x}\leq a^{x}+b^{x}$ for $x\in(0,1]$ and $a,b\geq 0$, and writing $w=\frac{3}{4}w_{1}+\frac{1}{2}w_{2}$, we obtain
$F_{X}\left(\sigma,H,T,w\right)\leq\left(\frac{\exp{\left(\kappa\left(1+\left(\frac{1}{4T}\left(1+\frac{2\delta}{\log{(hT)}}\right)\right)^{2}\right)\right)}}{1-H^{-1}}\right)^{\frac{1}{w}}\times\\ \times\left(\left(K_{1}^{\frac{\frac{3}{4}w_{1}}{w}}K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\right)T^{\frac{\frac{3}{4}w_{1}}{w}}\left(\log{T}\right)^{\frac{w_{1}}{w}}\left(\log{(hT)}\right)^{\frac{\frac{1}{2}w_{1}+w_{2}}{w}}+\left(K_{2}^{\frac{\frac{3}{4}w_{1}}{w}}K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\right)T^{\frac{\frac{3}{4}w_{1}}{w}}\left(\log{(hT)}\right)^{\frac{w_{2}}{w}}\right)$ (42)
by Corollary 6. We need to estimate the resulting terms on the right-hand side of (42). In what follows, remember that $\sigma\in[1/2,1]$. Because
$\frac{3(1-\sigma)}{2-\sigma}\leq\frac{\frac{3}{4}w_{1}}{w}=\frac{3(1-\sigma)+\frac{3\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{3(1-\sigma)}{2-\sigma}+\frac{3\delta}{\log{(hT_{0})}},$
we have
$K_{1}^{\frac{\frac{3}{4}w_{1}}{w}}\leq\max\left\{1,K_{1}^{\frac{3\delta}{\log{(hT_{0})}}}\right\}K_{1}^{\frac{3(1-\sigma)}{2-\sigma}}$ (43)
and
$T^{\frac{\frac{3}{4}w_{1}}{w}}\leq\exp{\left(\frac{3\delta\log{T}}{\log{(hT)}}\right)}T^{\frac{3(1-\sigma)}{2-\sigma}}\leq\exp{\left(3\delta\max\left\{1,\frac{\log{T_{0}}}{\log{(hT_{0})}}\right\}\right)}T^{\frac{3(1-\sigma)}{2-\sigma}}.$ (44)
Inequality (43) also holds for $K_{2}$ in place of $K_{1}$. Furthermore, because
$\frac{2\sigma-1}{2-\sigma}-\frac{3\delta}{\log{(hT_{0})}}\leq\frac{\frac{1}{2}w_{2}}{w}=\frac{2\sigma-1}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{2\sigma-1}{2-\sigma},$
we also have
$K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\leq\max\left\{1,K_{3}^{-\frac{3\delta}{\log{(hT_{0})}}}\right\}K_{3}^{\frac{2\sigma-1}{2-\sigma}}.$ (45)
Turning to the log-terms, because $\log{T}\geq 1$ and
$\frac{w_{1}}{w}=\frac{4(1-\sigma)+\frac{4\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{4(1-\sigma)}{2-\sigma}+\frac{4\delta}{\log{(hT)}},$
we have
$\left(\log{T}\right)^{\frac{w_{1}}{w}}\leq\exp{\left(\frac{4\delta\log{\log{T}}}{\log{(hT)}}\right)}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\leq\exp{\left(4\delta\max\left\{\frac{\log{\log{T_{0}}}}{\log{(hT_{0})}},\frac{\log{\log{T_{0}}}}{\log{T_{0}}}\right\}\right)}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}},$
where we also used $T_{0}\geq e^{e}$. Moreover,
$\frac{\frac{1}{2}w_{1}+w_{2}}{w}=\frac{2\sigma+\frac{2\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{2\sigma}{2-\sigma},\quad\frac{w_{2}}{w}=\frac{2(2\sigma-1)}{2-\sigma+\frac{3\delta}{\log(hT)}}\leq\frac{2(2\sigma-1)}{2-\sigma},$
which implies
$\left(\log{(hT)}\right)^{\frac{\frac{1}{2}w_{1}+w_{2}}{w}}\leq\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}},\quad\left(\log{(hT)}\right)^{\frac{w_{2}}{w}}\leq\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}$ (46)
since $\log{(hT)}\geq 1$. Finally,
$\frac{1}{w}=\frac{2+\frac{4\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{2}{2-\sigma}$
is used to bound the first term on the right-hand side of (42), and this contributes $L_{0}$ from (35) to the final estimate. Inequalities (43)–(46) contribute to $L_{1}$ from (39), while inequality (43) for $K_{2}$ in place of $K_{1}$, and inequalities (44), (45) and (46), contribute to $L_{2}$ from (40). The proof of Lemma 9 is thus complete. ∎
Proof of Theorem 1.
We are using (32) with $\sigma^{\prime}=\sigma-d/\log{T}$ and $\sigma^{\prime\prime}=3/2$. Note that $\sigma^{\prime}\in(1/2,1)$. Temporarily assume the non-vanishing condition on $h_{X}(s)$. The second integral in (32) is estimated by Lemma 7, while the last two integrals are estimated by Lemma 8. Note that the conditions of the lemmas are satisfied by the conditions of Theorem 1. Therefore, it remains to estimate the first integral in (32). Because $\log{\left(1+u^{2}\right)}\leq u^{a}$ for $u\geq 0$ and $a\in[4/3,2]$, we have
$\log{\left|h_{X}(s)\right|}\leq\log{\left(1+\left|f_{X}(s)\right|^{2}\right)}\leq\left|f_{X}(s)\right|^{\frac{1}{w}},\quad w=\frac{3}{4}w_{1}\left(\sigma^{\prime},\frac{1}{2},1+\frac{\delta}{\log{X}}\right)+\frac{1}{2}w_{2}\left(\sigma^{\prime},\frac{1}{2},1+\frac{\delta}{\log{X}}\right)$
for $\delta>0$, and with $w_{1}$ and $w_{2}$ as in (21). Thus,
$\int_{H}^{T}\log{\left|h_{X}\left(\sigma^{\prime}+\mathrm{i}t\right)\right|}\mathrm{d}t\leq F_{X}\left(\sigma^{\prime},H,T,w\right).$
Taking $X=hT$, we can see that all other conditions of Lemma 9 are satisfied by the conditions of Theorem 1. Therefore,
$F_{X}$ $\left(\sigma^{\prime},H,T,w\right)\leq\max\left\{1,L_{0}^{-\frac{2d}{\log{T_{0}}}}\right\}\max\left\{1,K_{3}^{-\frac{3d}{\log{T_{0}}}}\right\}L_{0}^{\frac{2}{2-\sigma}}K_{3}^{\frac{2\sigma-1}{2-\sigma}}\times$
×$\times\left(L_{1}\max\left\{1,K_{1}^{\frac{3d}{\log{T_{0}}}}\right\}\exp{\left(d\left(3+\frac{4\log{\log{T_{0}}}}{\log{T_{0}}}\right)\right)}\left(K_{1}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}}\right.$
$\left.+L_{2}e^{3d}\max\left\{1,K_{2}^{\frac{3d}{\log{T_{0}}}}\right\}\left(K_{2}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}\right),$
where $L_{0}$, $L_{1}$, $L_{2}$ and $K_{1}$, $K_{2}$, $K_{3}$ are as in Lemma 9. It is not hard to see that the statement of Theorem 1 now follows after combining all estimates. Also, the non-vanishing condition can be removed by evoking the standard perturbation argument since the right-hand side of the main estimate of Theorem 1 is a continuous function in $\sigma$, $T$ and $H$. ∎
5. The fourth power moment of $\zeta(s)$ on the critical line
In this section, we prove Theorem 2. The proof, which is presented through several subsections in Section 5.3, is technically quite involved and requires estimates on various sums that include the divisor function $d(n)$ and its square. These results, which might be of interest in their own right, are gathered in Section 5.2. Finally, we present the method together with the plan of the proof of Theorem 2 in the following section.
5.1. The method
Ingham’s proof of (7) is based on an approximate functional equation for $\zeta^{2}(s)$ in symmetric form and is relatively complicated. A shorter proof based on a Tauberian theorem for integrals but that produces only an asymptotic estimate (without an error term) can be found in [Tit86, Section 7]. However, a very elegant proof of (7) was found by Ramachandra [Ram75]. His proof, which is reproduced and extended in [Ivi03, Section 5.3], is based on a qualitative version of the following mean value theorem for Dirichlet polynomials.
Lemma 10.
Let $T_{2}\geq T_{1}>0$ and let $\{a_{n}\}_{n=1}^{\infty}$ be a sequence of complex numbers. Then
$\left|\int_{T_{1}}^{T_{2}}\left|\sum_{n=1}^{\infty}a_{n}n^{\mathrm{i}t}\right|^{2}\mathrm{d}t-\left(T_{2}-T_{1}\right)\sum_{n=1}^{\infty}|a_{n}|^{2}\right|\leq 2\pi m_{0}\sum_{n=1}^{\infty}\left(n+\frac{1}{2}\right)|a_{n}|^{2},$ (47)
where $m_{0}$ is from Lemma 5, provided that $\sum_{n=1}^{\infty}|a_{n}|\left(1+n|a_{n}|\right)$ converges.
Proof.
Let $X\geq 1$. We have
$\int_{T_{1}}^{T_{2}}\left|\sum_{n\leq X}a_{n}n^{\mathrm{i}t}\right|^{2}\mathrm{d}t=\left(T_{2}-T_{1}\right)\sum_{n\leq X}|a_{n}|^{2}-\mathrm{i}\mathop{\sum\sum}_{1\leq n\neq m\leq X}a_{n}\overline{a_{m}}\frac{(n/m)^{\mathrm{i}T_{2}}-(n/m)^{\mathrm{i}T_{1}}}{\log{n}-\log{m}}.$
Inequality (47) now follows after employing [Pre84, Inequality (D)] twice on the above double sum, noticing that $\left(\log{\frac{n+1}{n}}\right)^{-1}\leq n+\frac{1}{2}$ for all $n\in\mathbb{N}$, and then taking $X\to\infty$ since all series are convergent due to the assumption. ∎
The functional equation for $\zeta(s)$ can be written as $\zeta(s)=\chi(s)\zeta(1-s)$, where
$\chi(s)\mathrel{\mathop{:}}=\frac{(2\pi)^{s}}{2\Gamma(s)\cos{\left(\frac{\pi s}{2}\right)}}.$ (48)
This is a meromorphic function on $\mathbb{C}$ with only simple poles at $s=2k-1$ for $k\in\mathbb{N}$. Also note that the functional equation implies $\chi(1/2+\mathrm{i}t)\chi(1/2-\mathrm{i}t)=\left|\chi(1/2\pm\mathrm{i}t)\right|^{2}=1$. The starting point in Ramachandra’s proof (which itself resembles the reflection principle [Ivi03, Section 4.4] or a half-way derivation of a weighted approximate functional equation [Ivi13, Section 4.5]) is to use the Mellin integral for $e^{-x}$ to get
$\sum_{n=1}^{\infty}\frac{d(n)}{n^{1/2+\mathrm{i}t}}e^{-n/T}=\frac{1}{2\pi\mathrm{i}}\int_{2-\mathrm{i}\infty}^{2+\mathrm{i}\infty}\zeta^{2}\left(\frac{1}{2}+\mathrm{i}t+z\right)\Gamma(z)T^{z}\mathrm{d}z$
for $T\geq 1$. By moving the line of integration to $\Re\{z\}=c_{1}\in(-1,-1/2)$ we get in addition to
$\frac{1}{2\pi\mathrm{i}}\int_{c_{1}-\mathrm{i}\infty}^{c_{1}+\mathrm{i}\infty}\zeta^{2}\left(\frac{1}{2}+\mathrm{i}t+z\right)\Gamma(z)T^{z}\mathrm{d}z$ (49)
also two terms that come from the simple pole of $\Gamma(z)$ at $z=0$ and double pole of $\zeta^{2}(1/2+\mathrm{i}t+z)$ at $z=1/2-\mathrm{i}t$. Next, the integral (49) can be written as
$\frac{1}{2\pi\mathrm{i}}\int_{c_{1}-\mathrm{i}\infty}^{c_{1}+\mathrm{i}\infty}\chi^{2}\left(\frac{1}{2}+\mathrm{i}t+z\right)\Gamma(z)T^{z}\left(\left(\sum_{n\leq T}+\sum_{n>T}\right)\frac{d(n)}{n^{1/2-\mathrm{i}t-z}}\right)\mathrm{d}z$ (50)
due to the functional equation. By moving the line of integration to $\Re\{z\}=c_{2}\in(0,1/2)$ only for the first integral (implied by the first sum) in (50) we get, beside the new integral, also another term that comes from the simple pole of $\Gamma(z)$ at $z=0$. After multiplying the final equation by $\chi^{-1}(1/2+\mathrm{i}t)$, and then truncating some sums and integrals, it is not hard to see that we obtain
$\chi^{-1}(1/2+\mathrm{i}t)\zeta^{2}(1/2+\mathrm{i}t)=\left|\zeta(1/2+\mathrm{i}t)\right|^{2}=\sum_{k=0}^{6}J_{k}(t),$ (51)
where
$J_{1}(t)\mathrel{\mathop{:}}=\chi{\left(\frac{1}{2}+\mathrm{i}t\right)}\sum_{n\leq T}\frac{d(n)}{n^{1/2-\mathrm{i}t}},\quad J_{2}(t)\mathrel{\mathop{:}}=\overline{J_{1}(t)},\quad J_{3}(t)\mathrel{\mathop{:}}=\chi{\left(\frac{1}{2}-\mathrm{i}t\right)}\sum_{n>T}\frac{d(n)}{n^{1/2+\mathrm{i}t}}e^{-n/T},$
$J_{4}(t)\mathrel{\mathop{:}}=\chi{\left(\frac{1}{2}-\mathrm{i}t\right)}\sum_{n\leq T}\frac{d(n)}{n^{1/2+\mathrm{i}t}}\left(e^{-n/T}-1\right),\quad J_{5}(t)\mathrel{\mathop{:}}=-\frac{1}{2\pi}\chi{\left(\frac{1}{2}-\mathrm{i}t\right)}\int_{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}P(0,1,c_{1};u)\mathrm{d}u,$
$J_{6}(t)\mathrel{\mathop{:}}=-\frac{1}{2\pi}\chi{\left(\frac{1}{2}-\mathrm{i}t\right)}\int_{-\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}}^{\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}}P(1,0,c_{2};u)\mathrm{d}u$
and
$J_{0}(t)\mathrel{\mathop{:}}=-\frac{1}{2\pi}\chi{\left(\frac{1}{2}-\mathrm{i}t\right)}\Biggl{(}\left(\int_{-\infty}^{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}+\int_{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\infty}\right)P(0,1,c_{1};u)\mathrm{d}u\\ \left.+\left(\int_{-\infty}^{-\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}}+\int_{\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}}^{\infty}\right)P(1,0,c_{2};u)\mathrm{d}u+2\pi\operatorname*{Res}_{z=\frac{1}{2}-\mathrm{i}t}\left(\zeta^{2}{\left(\frac{1}{2}+\mathrm{i}t+z\right)}\Gamma(z)T^{z}\right)\right)$
for $\mathfrak{a}_{1}>0$, $\mathfrak{a}_{2}>1$, $\mathfrak{b}_{1}>0$, $\mathfrak{b}_{2}>1$ and
$P(a,b,c;u)\mathrel{\mathop{:}}=\chi^{2}\left(\frac{1}{2}+c+\mathrm{i}(t+u)\right)\Gamma\left(c+\mathrm{i}u\right)T^{c+\mathrm{i}u}\left(\left(a\sum_{n\leq T}+b\sum_{n>T}\right)\frac{d(n)}{n^{\frac{1}{2}-c-\mathrm{i}(t+u)}}\right).$
Squaring (51), and then integrating the result from $T$ to $2T$ gives
$\mathcal{M}_{2}(T,2T)=\underbrace{2\int_{T}^{2T}\left|J_{1}(t)\right|^{2}\mathrm{d}t}_{\text{I. part}}+\underbrace{\int_{T}^{2T}\left(J_{1}^{2}(t)+J_{2}^{2}(t)\right)\mathrm{d}t}_{\text{II. part}}+\mathcal{R},$ (52)
where
$|\mathcal{R}|\leq\underbrace{2\sum_{k=3}^{6}\left|\int_{T}^{2T}\left(J_{1}(t)+J_{2}(t)\right)J_{k}(t)\mathrm{d}t\right|}_{\text{III. part}}+\underbrace{\sum_{k=3}^{6}\int_{T}^{2T}\left|J_{k}(t)\right|^{2}\mathrm{d}t}_{\text{IV. part}}\\ +\underbrace{2\mathop{\sum\sum}_{3\leq n<m\leq 6}\int_{T}^{2T}\left|J_{n}(t)J_{m}(t)\right|\mathrm{d}t}_{\text{V. part}}+\underbrace{2\int_{T}^{2T}\left|J_{0}(t)\right|\cdot\left|\zeta\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t+\int_{T}^{2T}\left|J_{0}(t)\right|^{2}\mathrm{d}t}_{\text{VI. part}}.$ (53)
The curly brackets in (52) and (53) indicate the plan of the proof of Theorem 2. Precisely, the seemingly unnatural appearance of the second and the third part allows us to obtain an asymptotically correct estimate (12) since we can show that all terms, except the first one, are $\ll T\log^{3}{T}$. However, if interested in obtaining just an upper bound like (13), then  (52) and (53) can be simplified to
$\mathcal{M}_{2}(T,2T)\leq 4\int_{T}^{2T}\left|J_{1}(t)\right|^{2}\mathrm{d}t+\sum_{k=3}^{6}\int_{T}^{2T}\left|J_{k}(t)\right|^{2}\mathrm{d}t+4\sum_{k=3}^{6}\int_{T}^{2T}\left|J_{1}(t)J_{k}(t)\right|\mathrm{d}t\\ +2\mathop{\sum\sum}_{3\leq n<m\leq 6}\int_{T}^{2T}\left|J_{n}(t)J_{m}(t)\right|\mathrm{d}t+2\int_{T}^{2T}\left|J_{0}(t)\right|\cdot\left|\zeta\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t+\int_{T}^{2T}\left|J_{0}(t)\right|^{2}\mathrm{d}t.$ (54)
Before proceeding to the proof, we need first some auxiliary results for various sums that contain $d(n)$ and $d^{2}(n)$, together with some estimates for the moduli of $\Gamma(z)$, $\chi(z)$ and $\zeta(1/2+\mathrm{i}t)$.
5.2. Some results on various sums that contain $d(n)$ and $d^{2}(n)$
In this section, we are going to prove several estimates on sums that involve $d(n)$ and $d^{2}(n)$, and are crucial for the estimation of terms in (52) and (53).
Lemma 11.
Let $x\geq x_{0}\geq 2$. Then
$\sum_{n\leq x}\frac{d(n)}{n^{\sigma}}\leq\frac{\mathcal{D}_{1}(x_{0},\sigma_{0})}{1-\sigma}x^{1-\sigma}\log{x},\quad\sigma_{0}\leq\sigma\leq 1/2,\;\sigma_{0}\leq 0,$
and
$\sum_{n>x}\frac{d(n)}{n^{\sigma}}\leq\frac{\mathcal{D}_{2}(x_{0},\sigma_{0})}{(\sigma_{0}-1)^{2}}x^{1-\sigma}\log{x},\quad\sigma\geq\sigma_{0}>1,$
where
$\mathcal{D}_{1}(x_{0},\sigma_{0})\mathrel{\mathop{:}}=1+\frac{2\gamma-1-\frac{\sigma_{0}}{1-\sigma_{0}}}{\log{x_{0}}}+\frac{(1-\sigma_{0})\left(\frac{1}{2}+|\sigma_{0}|-\sigma_{0}\right)}{\sqrt{x_{0}}}+\frac{1-2\gamma\sigma_{0}}{\sqrt{x_{0}}\log{x_{0}}},$ (55)
$\mathcal{D}_{2}(x_{0},\sigma_{0})\mathrel{\mathop{:}}=\sigma_{0}-1+\frac{1+2\gamma(\sigma_{0}-1)}{\log{x_{0}}}+\frac{4\sigma_{0}-1}{(2\sigma_{0}-1)\sqrt{x_{0}}\log{x_{0}}}$ (56)
and $\gamma$ is the Euler–Mascheroni constant.
Proof.
By [BBR12, Theorem 1.1] we have $\sum_{n\leq x}d(n)=x\log{x}+(2\gamma-1)x+\Delta(x)$ with $\left|\Delta(x)\right|\leq\sqrt{x}$ for $x\geq 1$. Let $1\leq X\leq Y$. By partial summation,
$\sum_{X<n\leq Y}\frac{d(n)}{n^{\sigma}}=Y^{-\sigma}\sum_{n\leq Y}d(n)-X^{-\sigma}\sum_{n\leq X}d(n)+\sigma\int_{X}^{Y}\frac{\sum_{n\leq u}d(n)}{u^{1+\sigma}}\mathrm{d}u.$ (57)
The statement of Lemma 11 for $\sigma_{0}\leq\sigma\leq 1/2$, $\sigma_{0}\leq 0$ now follows by taking $X=1$ and $Y=x$ in (57), by straightforward integration, by noticing that
$\frac{\sigma\left(1-x^{1-\sigma}\right)}{(1-\sigma)^{2}}\leq\left\{\begin{array}[]{ll}0,&\sigma\in[0,1),\\ \frac{-\sigma x^{1-\sigma}}{(1-\sigma)^{2}},&\sigma<0,\end{array}\right.$
and by the mean-value estimate
$\frac{\left(1+2\left(|\sigma|-\sigma\right)\right)x^{1/2-\sigma}-2|\sigma|}{1-2\sigma}\leq 1+\left(\frac{1}{2}+|\sigma|-\sigma\right)x^{\frac{1}{2}-\sigma}\log{x}$
that is valid for all $\sigma<1/2$. The corresponding result for $\sigma\geq\sigma_{0}>1$ simply follows by taking $X=x$ and $Y\to\infty$ in (57), while observing that $(\sigma-1)^{-2}\mathcal{D}_{2}(x_{0},\sigma)$ is a decreasing function in $\sigma>1$. ∎
Lemma 12.
Let $x\geq x_{0}\geq 2$ and $\sigma\in[0,1]$. Then
$\sum_{n\leq x}\frac{d(n)}{n^{\sigma}}\left(1-e^{-\frac{n}{x}}\right)\leq\frac{\mathcal{D}_{1}(x_{0},\sigma-1)}{2-\sigma}x^{1-\sigma}\log{x},\quad\sum_{n>x}\frac{d(n)}{n^{\sigma}}e^{-\frac{n}{x}}\leq\mathcal{D}_{3}(x_{0})x^{1-\sigma}\log{x},$
where $\mathcal{D}_{1}(x_{0},\sigma_{0})$ is defined by (55) and
$\mathcal{D}_{3}(x_{0})\mathrel{\mathop{:}}=\frac{1}{e}+\frac{1+2\gamma}{e\log{x_{0}}}+\frac{5}{2e\sqrt{x_{0}}\log{x_{0}}},$ (58)
where $\gamma$ is the Euler–Mascheroni constant.
Proof.
The first inequality simply follows from Lemma 11 due to the fact that $\left|1-e^{-u}\right|\leq u$ for $u\in[0,1]$ and $\sigma-1\in[-1,0]$. Concerning the second inequality, we have
$\sum_{n>x}\frac{d(n)}{n^{\sigma}}e^{-\frac{n}{x}}\leq\frac{1}{x^{\sigma}}\sum_{n>x}d(n)e^{-\frac{n}{x}},\quad\sigma\geq 0,$
and
$\sum_{n>x}d(n)e^{-\frac{n}{x}}=-\frac{1}{e}\sum_{n\leq x}d(n)+\frac{1}{x}\int_{x}^{\infty}e^{-\frac{u}{x}}\sum_{n\leq u}d(n)\mathrm{d}u$
by partial summation. Now,
$\frac{1}{x}\int_{x}^{\infty}e^{-\frac{u}{x}}\sum_{n\leq u}d(n)\mathrm{d}u$ $=\frac{1}{x}\int_{x}^{\infty}e^{-\frac{u}{x}}\left(u\log{u}+(2\gamma-1)u\right)\mathrm{d}u+\frac{1}{x}\int_{x}^{\infty}e^{-\frac{u}{x}}\Delta(u)\mathrm{d}u$
$\leq\frac{2}{e}x\log{x}+\frac{4\gamma}{e}x+\frac{3}{2e}\sqrt{x},$
where we have used
$\int_{x}^{\infty}e^{-\frac{u}{x}}u^{\alpha}(\log{u})^{k}\mathrm{d}u=\frac{1+\alpha}{e}x^{1+\alpha}(\log{x})^{k}+\alpha(\alpha-1)x^{2}\int_{x}^{\infty}e^{-\frac{u}{x}}u^{\alpha-2}(\log{u})^{k}\mathrm{d}u\\ +\alpha kx^{2}\int_{x}^{\infty}e^{-\frac{u}{x}}u^{\alpha-2}(\log{u})^{k-1}\mathrm{d}u+kx\int_{x}^{\infty}e^{-\frac{u}{x}}u^{\alpha-1}(\log{u})^{k-1}\mathrm{d}u,$
which is valid for all $\alpha\in\mathbb{R}$ and $k\in\mathbb{R}$. The stated result now easily follows. ∎
Lemma 13.
Let $x\geq 2$. Then
$\sum_{n\leq x}d^{2}(n)=D_{1}x\log^{3}{x}+D_{2}x\log^{2}{x}+D_{3}x\log{x}+D_{4}x+D_{0}(x),$
where $D_{1}=1/\pi^{2}$, $0.744341<D_{2}<0.744342$, $0.823265<D_{3}<0.823266$, $0.460323<D_{4}<0.460324$ and $|D_{0}(x)|\leq 9.73x^{3/4}\log{x}$.
Proof.
This is [CHT21, Theorem 2]. ∎
Lemma 14.
Let $x\geq x_{0}\geq 2$. Then
$\sum_{n\leq x}\frac{d^{2}{(n)}}{n^{\sigma}}\leq\mathcal{D}_{4}(x_{0},\sigma)x^{1-\sigma}\log^{3}{x},\quad\sigma\in[-3,1),$
and
$\sum_{n>x}\frac{d^{2}{(n)}}{n^{\sigma}}\leq\frac{\mathcal{D}_{5}(x_{0},\sigma)}{(\sigma-1)^{4}}x^{1-\sigma}\log^{3}{x},\quad\sigma>1,$
where
$\mathcal{D}_{4}(x_{0},\sigma)\mathrel{\mathop{:}}=\left\{\begin{array}[]{ll}\mathcal{D}_{41}(x_{0},\sigma),&\sigma\in[-3,0),\\ \frac{1}{1-\sigma}\mathcal{D}_{42}(x_{0},\sigma),&\sigma\in[0,1)\end{array}\right.$ (59)
with
$\mathcal{D}_{41}(x_{0},\sigma)$ $\mathrel{\mathop{:}}=\frac{1}{(1-\sigma)\pi^{2}}+\left(-\frac{3\sigma}{(1-\sigma)^{2}\pi^{2}}+\frac{0.744342}{1-\sigma}\right)\frac{1}{\log{x_{0}}}$
$+\left(\frac{6\sigma}{(1-\sigma)^{3}\pi^{2}}-\frac{1.488684\sigma}{(1-\sigma)^{2}}+\frac{0.823266}{1-\sigma}\right)\frac{1}{\log^{2}{x_{0}}}$
$+\left(-\frac{6\sigma}{(1-\sigma)^{4}\pi^{2}}+\frac{1.488684\sigma}{(1-\sigma)^{3}}-\frac{0.823266\sigma}{(1-\sigma)^{2}}+\frac{0.460324}{1-\sigma}\right)\frac{1}{\log^{3}{x_{0}}}+\frac{17.52}{x_{0}^{1/4}\log^{2}{x_{0}}},$
and
$\mathcal{D}_{42}(x_{0},\sigma)\mathrel{\mathop{:}}=\frac{1}{\pi^{2}}+\frac{0.744342}{\log{x_{0}}}+\frac{0.823266}{\log^{2}{x_{0}}}+\frac{0.460324}{\log^{3}{x_{0}}}+\frac{9.73\sigma(1-\sigma)}{x_{0}^{\min\{1/4,1-\sigma\}}\log{x_{0}}}+\frac{9.73(1-\sigma)}{x_{0}^{1/4}\log^{2}{x_{0}}},$
while
$\mathcal{D}_{5}(x_{0},\sigma)$ $\mathrel{\mathop{:}}=\frac{(\sigma-1)^{3}}{\pi^{2}}+\left(\sigma-1\right)^{2}\left(\frac{3\sigma}{\pi^{2}}+0.744342(\sigma-1)\right)\frac{1}{\log{x_{0}}}$
$+\left(\sigma-1\right)\left(\frac{6\sigma}{\pi^{2}}+1.488684\sigma(\sigma-1)+0.823266(\sigma-1)^{2}\right)\frac{1}{\log^{2}{x_{0}}}$
$+\left(\frac{6\sigma}{\pi^{2}}+1.488684\sigma(\sigma-1)+0.823266\sigma(\sigma-1)^{2}+0.460324(\sigma-1)^{3}\right)\frac{1}{\log^{3}{x_{0}}}$
$+\left(1+\frac{\sigma}{\sigma-3/4}\left(1+\frac{1}{\left(\sigma-3/4\right)\log{x_{0}}}\right)\right)\frac{9.73(\sigma-1)^{4}}{x_{0}^{1/4}\log^{2}{x_{0}}}.$ (60)
Proof.
Let $2\leq X\leq Y$, $\sigma\in\mathbb{R}\setminus\{1\}$ and $k\geq 0$. Note that
$\int_{X}^{Y}u^{-\sigma}\left(\log{u}\right)^{k}\mathrm{d}u=\frac{\log^{k}{Y}}{(1-\sigma)Y^{\sigma-1}}-\frac{\log^{k}{X}}{(1-\sigma)X^{\sigma-1}}-\frac{k}{1-\sigma}\int_{X}^{Y}u^{-\sigma}\left(\log{u}\right)^{k-1}\mathrm{d}u.$
Repeated use of the above recursive relation, in combination with partial summation and Lemma 13, gives
$\sum_{X<n\leq Y}\frac{d^{2}{(n)}}{n^{\sigma}}$ $=Y^{-\sigma}\sum_{n\leq Y}d^{2}{(n)}-X^{-\sigma}\sum_{n\leq X}d^{2}{(n)}+\sigma\int_{X}^{Y}\frac{\sum_{n\leq u}d^{2}{(n)}}{u^{1+\sigma}}\mathrm{d}u$
$=\sum_{k=1}^{4}D_{k}Y^{1-\sigma}\left(\log{Y}\right)^{4-k}+\sigma\mathcal{Z}(Y,\sigma)-\sum_{k=1}^{4}D_{k}X^{1-\sigma}\left(\log{X}\right)^{4-k}-\sigma\mathcal{Z}(X,\sigma)$
$+Y^{-\sigma}D_{0}(Y)-X^{-\sigma}D_{0}(X)+\sigma\int_{X}^{Y}\frac{D_{0}(u)}{u^{1+\sigma}}\mathrm{d}u,$ (61)
where
$\mathcal{Z}(Z,\sigma)$ $\mathrel{\mathop{:}}=\frac{D_{1}\log^{3}{Z}}{(1-\sigma)Z^{\sigma-1}}+\left(-\frac{3D_{1}}{(1-\sigma)^{2}}+\frac{D_{2}}{1-\sigma}\right)\frac{\log^{2}{Z}}{Z^{\sigma-1}}+\left(\frac{6D_{1}}{(1-\sigma)^{3}}-\frac{2D_{2}}{(1-\sigma)^{2}}+\frac{D_{3}}{1-\sigma}\right)\frac{\log{Z}}{Z^{\sigma-1}}$
$+\left(-\frac{6D_{1}}{(1-\sigma)^{4}}+\frac{2D_{2}}{(1-\sigma)^{3}}-\frac{D_{3}}{(1-\sigma)^{2}}+\frac{D_{4}}{1-\sigma}\right)\frac{1}{Z^{\sigma-1}}.$ (62)
Also,
$\mathcal{Z}(Y,\sigma)-\mathcal{Z}(X,\sigma)\leq\left(D_{1}\log^{3}{Y}+D_{2}\log^{2}{Y}+D_{3}\log{Y}+D_{4}\right)\frac{Y^{1-\sigma}-X^{1-\sigma}}{1-\sigma}.$ (63)
Let $\sigma\in[-3,0]$, $X=2$ and $Y=x$. Then (5.2) and (5.2) imply
$\sum_{n\leq x}\frac{d^{2}{(n)}}{n^{\sigma}}$ $\leq\left(\frac{D_{1}}{1-\sigma}+\left(-\frac{3\sigma D_{1}}{(1-\sigma)^{2}}+\frac{D_{2}}{1-\sigma}\right)\frac{1}{\log{x}}+\left(\frac{6\sigma D_{1}}{(1-\sigma)^{3}}-\frac{2\sigma D_{2}}{(1-\sigma)^{2}}+\frac{D_{3}}{1-\sigma}\right)\frac{1}{\log^{2}{x}}\right.$
$\left.+\left(-\frac{6\sigma D_{1}}{(1-\sigma)^{4}}+\frac{2\sigma D_{2}}{(1-\sigma)^{3}}-\frac{\sigma D_{3}}{(1-\sigma)^{2}}+\frac{D_{4}}{1-\sigma}\right)\frac{1}{\log^{3}{x}}\right)x^{1-\sigma}\log^{3}{x}$
$+1-2^{-\sigma}-\sigma\mathcal{Z}(2,\sigma)+9.73x^{\frac{3}{4}-\sigma}\log{x}+9.73|\sigma|\int_{2}^{x}\frac{\log{u}}{u^{1/4+\sigma}}\mathrm{d}u.$ (64)
Note that
$\int_{2}^{x}\frac{\log{u}}{u^{1/4+\sigma}}\mathrm{d}u\leq\frac{x^{3/4-\sigma}-2^{3/4-\sigma}}{3/4-\sigma}\log{x}\leq x^{\max\{3/4-\sigma,0\}}\log^{2}{x}-2^{-1/4}\log{2}\log{x}$ (65)
for $\sigma\in(-\infty,1]$ by the mean-value theorem. The values of the expressions in the inner brackets of (5.2) are positive, while the last line in (5.2) is $\leq 17.52x^{3/4-\sigma}\log{x}$ by the first inequality in (65). The corresponding result of Lemma 14 now easily follows. Next, consider $\sigma\in[0,1)$, $X=2$ and $Y=x$. It follows from (5.2), (63) and (65) that
$\sum_{n\leq x}\frac{d^{2}(n)}{n^{\sigma}}\leq\frac{x^{1-\sigma}}{1-\sigma}\sum_{k=1}^{4}D_{k}\left(\log{x}\right)^{4-k}+1-2^{-\sigma}-\frac{\sigma}{1-\sigma}\sum_{k=1}^{4}D_{k}\left(\log{2}\right)^{4-k}\\ +9.73x^{\frac{3}{4}-\sigma}\log{x}+9.73\sigma x^{\max\{3/4-\sigma,0\}}\log^{2}{x}-9.73\sigma 2^{-1/4}\log^{2}{2}.$
The sum of those terms on the right-hand side of the above inequality that do not depend on $x$ is not positive. The corresponding result of Lemma 14 now easily follows. Finally, let $\sigma>1$, $X=x$ and $Y\to\infty$. The result follows from (5.2) and (5.2), together with
$\int_{x}^{\infty}\frac{\log{u}}{u^{1/4+\sigma}}\mathrm{d}u=\frac{1}{(\sigma-3/4)x^{\sigma-3/4}}\left(\log{x}+\frac{1}{\sigma-3/4}\right).$
The proof is thus complete. ∎
Lemma 15.
Let $x\geq x_{0}\geq 2$. Then
$\left|\sum_{n\leq x}\frac{d^{2}(n)}{n}-\frac{1}{4\pi^{2}}\log^{4}{x}\right|\leq\mathcal{D}_{6}(x_{0})\log^{3}{x},$
where
$\mathcal{D}_{6}(x_{0})\mathrel{\mathop{:}}=0.349436+\frac{1.155975}{\log{x_{0}}}+\frac{1.28359}{\log^{2}{x_{0}}}+\frac{154}{\log^{3}{x_{0}}}+\frac{9.73}{x_{0}^{1/4}\log^{2}{x_{0}}}.$ (66)
Proof.
The result follows by straightforward integration, using the first equality in (5.2) for $\sigma=1$, $X=2$ and $Y=x$ in combination with Lemma 13. ∎
Lemma 16.
Let $x\geq x_{0}\geq e^{4}$ and $\sigma\geq-1$. Then
$\sum_{n>x}\frac{d^{2}(n)}{n^{\sigma}}e^{-\frac{2n}{x}}\leq\mathcal{D}_{7}(x_{0})x^{1-\sigma}\log^{3}{x},$
where
$\mathcal{D}_{7}(x_{0})\mathrel{\mathop{:}}=\frac{3}{(2e\pi)^{2}}+\frac{0.12441}{\log{x_{0}}}+\frac{0.364}{\log^{2}{x_{0}}}+\frac{0.3}{\log^{3}{x_{0}}}+\frac{4.61}{x_{0}^{1/4}\log^{2}{x_{0}}}.$ (67)
Proof.
We are going to show that
$\sum_{n>x}nd^{2}(n)e^{-\frac{2n}{x}}\leq\mathcal{D}_{7}(x_{0})x^{2}\log^{3}{x}.$ (68)
Then the result will follow since
$\sum_{n>x}\frac{d^{2}(n)}{n^{\sigma}}e^{-\frac{2n}{x}}\leq\frac{1}{x^{1+\sigma}}\sum_{n>x}nd^{2}(n)e^{-\frac{2n}{x}}$
for $\sigma\geq-1$. We have
$\sum_{n>x}nd^{2}(n)e^{-\frac{2n}{x}}$ $=-e^{-2}x\sum_{n\leq x}d^{2}(n)+\int_{x}^{\infty}\left(\frac{2u}{x}-1\right)e^{-\frac{2u}{x}}\sum_{n\leq u}d^{2}(n)\mathrm{d}u$
$=-e^{-2}x^{2}\sum_{k=1}^{4}D_{k}(\log{x})^{4-k}+\sum_{k=1}^{4}D_{k}\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}(\log{u})^{4-k}\mathrm{d}u$
$-e^{-2}xD_{0}(x)+\int_{x}^{\infty}\left(\frac{2u}{x}-1\right)e^{-\frac{2u}{x}}D_{0}(u)\mathrm{d}u$ (69)
by partial summation and Lemma 13. Repeated use of partial integration gives
$\int_{x}^{\infty}e^{-\frac{2u}{x}}$ $u^{\alpha}(\log{u})^{k}\mathrm{d}u=\frac{\alpha^{2}+\alpha+4}{8e^{2}}x^{1+\alpha}(\log{x})^{k}+\frac{k(2\alpha+1)}{8e^{2}}x^{1+\alpha}(\log{x})^{k-1}+\frac{k(k-1)}{8e^{2}}x^{1+\alpha}(\log{x})^{k-2}$
$+\frac{\alpha(\alpha-1)(\alpha-2)}{8}x^{3}\int_{x}^{\infty}e^{-\frac{2u}{x}}\frac{(\log{u})^{k}}{u^{3-\alpha}}\mathrm{d}u+\frac{k\left(3\alpha^{2}-6\alpha+2\right)}{8}x^{3}\int_{x}^{\infty}e^{-\frac{2u}{x}}\frac{(\log{u})^{k-1}}{u^{3-\alpha}}\mathrm{d}u$
$+\frac{3(\alpha-1)k(k-1)}{8}x^{3}\int_{x}^{\infty}e^{-\frac{2u}{x}}\frac{(\log{u})^{k-2}}{u^{3-\alpha}}\mathrm{d}u+\frac{k(k-1)(k-2)}{8}x^{3}\int_{x}^{\infty}e^{-\frac{2u}{x}}\frac{(\log{u})^{k-3}}{u^{3-\alpha}}\mathrm{d}u$
for any $\alpha\in\mathbb{R}$ and $k\in\mathbb{R}$. Then
$-e^{-2}xD_{0}(x)+\int_{x}^{\infty}\left(\frac{2u}{x}-1\right)e^{-\frac{2u}{x}}D_{0}(u)\mathrm{d}u$ $\leq\frac{9.73}{e^{2}}x^{\frac{7}{4}}\log{x}+\frac{19.46\log{x}}{x^{5/4}}\int_{x}^{\infty}e^{-\frac{2u}{x}}u^{2}\mathrm{d}u$
$\leq 4.61x^{\frac{7}{4}}\log{x}.$ (70)
Using
$\int_{x}^{\infty}e^{-\frac{2u}{x}}\frac{(\log{u})^{k}}{u^{\alpha}}\mathrm{d}u\leq\frac{1}{2e^{2}}x^{1-\alpha}(\log{x})^{k},\quad k\leq 2,\alpha\geq 1,$
we also obtain
$\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}(\log{u})^{3}\mathrm{d}u\leq\frac{7}{4e^{2}}x^{2}(\log{x})^{3}+\frac{57}{16e^{2}}x^{2}(\log{x})^{2}+\frac{3}{e^{2}}x^{2}\log{x}+\frac{3}{4e^{2}}x^{2},$
$\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}(\log{u})^{2}\mathrm{d}u\leq\frac{7}{4e^{2}}x^{2}(\log{x})^{2}+\frac{19}{8e^{2}}x^{2}\log{x}+\frac{1}{e^{2}}x^{2},$
$\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}\log{u}\mathrm{d}u\leq\frac{7}{4e^{2}}x^{2}\log{x}+\frac{19}{16e^{2}}x^{2},$
$\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}\mathrm{d}u\leq\frac{7}{4e^{2}}x^{2}.$
Therefore,
$-e^{-2}x^{2}\sum_{k=1}^{4}D_{k}(\log{x})^{4-k}+\sum_{k=1}^{4}D_{k}\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}(\log{u})^{4-k}\mathrm{d}u\leq\frac{3D_{1}}{4e^{2}}x^{2}(\log{x})^{3}\\ +\left(\frac{57D_{1}}{16e^{2}}+\frac{3D_{2}}{4e^{2}}\right)x^{2}(\log{x})^{2}+\left(\frac{3D_{1}}{e^{2}}+\frac{19D_{2}}{8e^{2}}+\frac{3D_{3}}{4e^{2}}\right)x^{2}\log{x}+\left(\frac{3D_{1}}{4e^{2}}+\frac{D_{2}}{e^{2}}+\frac{19D_{3}}{16e^{2}}+\frac{3D_{4}}{4e^{2}}\right)x^{2}.$ (71)
It is straightforward to verify that (68) is valid by considering (5.2) and (71) in (5.2). ∎
Lemma 17.
Let $x\geq x_{0}\geq 2$. Then
$\sum_{n\leq x}\frac{d^{2}(n)}{n^{\sigma}}\left(e^{-\frac{n}{x}}-1\right)^{2}\leq\mathcal{D}_{4}(x_{0},\sigma-2)x^{1-\sigma}\log^{3}{x}$
for $\sigma\in[-1,2]$, where $\mathcal{D}_{4}(x_{0},\sigma)$ is defined by (59).
Proof.
We have
$\sum_{n\leq x}\frac{d^{2}(n)}{n^{\sigma}}\left(e^{-\frac{n}{x}}-1\right)^{2}\leq\frac{1}{x^{2}}\sum_{n\leq x}d^{2}(n)n^{2-\sigma}$
since $\left|e^{-u}-1\right|\leq u$ for $u\in[0,1]$. The result now follows by Lemma 14. ∎
Remark 2.
The main terms in the estimates from Lemmas 1216 and 17 can be improved for certain $\sigma$’s, e.g., essentially the same method as in the proof of Lemma 16 gives
$\sum_{n>x}d^{2}(n)e^{-\frac{2n}{x}}=\frac{1}{2(e\pi)^{2}}x\log^{3}{x}+O\left(x\log^{2}{x}\right).$
However, there are some technical difficulties to obtain asymptotically correct estimates when $\sigma$ is not an integer. Note that the above also shows that the first estimate of [Ivi03, p. 138] is not correct, although this does not change the final result.
5.3. Proof of Theorem 2
In this section, we provide the proof of Theorem 2 by estimating each of the six parts in (52) and (53). In order to simplify some expressions throughout the following sections, we define
$\eta(T_{0})\mathrel{\mathop{:}}=1+\frac{\pi m_{0}}{T_{0}},$
where $m_{0}$ is from Lemma 5. In addition, we assume the notation and conditions (1)–(6) from Theorem 2. Before proceeding to the proof, we need two additional results that are related to the Gamma function.
Lemma 18.
Let $z=x+\mathrm{i}y$ with $x\in[-1,1]$ and $|y|\geq y_{0}\geq 2$. Then
$\exp{\left(-\frac{1}{2y_{0}^{2}}\right)}\leq\frac{\left|\Gamma(z)\right|}{\sqrt{2\pi}|z|^{x-\frac{1}{2}}e^{-\frac{\pi}{2}|y|}}\leq\exp{\left(\frac{1}{2y_{0}^{2}}\right)}$ (72)
and
$\mathcal{U}^{-}(y_{0})\leq\left(\frac{|z|}{2\pi}\right)^{x-\frac{1}{2}}\left|\chi(z)\right|\leq\mathcal{U}^{+}(y_{0}),\quad\mathcal{U}^{\pm}(y_{0})\mathrel{\mathop{:}}=\left(1\mp e^{-\pi y_{0}}\right)^{-1}\exp{\left(\pm\frac{1}{2y_{0}^{2}}\right)}.$ (73)
In addition, $|\chi(z)|\leq 0.4|y|^{1/2-x}+16$ for $x\in[-1/2,0]$ and $y\in\mathbb{R}$, while $|\chi(z)|\leq 6.31/(1-x)$ for $x\in[1/2,1)$ and $y\in\mathbb{R}$.
Proof.
Inequalities (72) follow from [DZA22, Theorem 2.1 (B2)] with $\theta=\arctan{2}$. Inequalities (73) follow from (48), (72) and
$1-e^{-\pi y_{0}}\leq 2e^{-\frac{\pi}{2}|y|}\left|\cos{\left(\frac{\pi z}{2}\right)}\right|\leq 1+e^{-\pi y_{0}},$
with the last inequalities true by [DZA22, Proposition 2.2]. For the second part, one can use (73) to obtain $|\chi(z)|\leq 0.4|y|^{1/2-x}$ for $x\in[-1/2,0]$ and $|y|\geq 10^{2}$, while the maximum principle and numerical verification show that $|\chi(z)|\leq 16$ holds for $x\in[-1/2,0]$ and $|y|\leq 10^{2}$. Finally, $|\chi(z)|\leq 1.1$ for $x\in[1/2,1]$ and $|y|\geq 2\pi$, while $|(z-1)\chi(z)|\leq 6.31$ for $x\in[1/2,1]$ and $|y|\leq 2\pi$ by the maximum principle and numerical verification. ∎
Lemma 19.
Let $t\geq T\geq 5$. Then
$\left|\operatorname*{Res}_{z=\frac{1}{2}-\mathrm{i}t}\left(\zeta^{2}{\left(\frac{1}{2}+\mathrm{i}t+z\right)}\Gamma(z)T^{z}\right)\right|\leq 12\sqrt{T}e^{-\frac{\pi}{2}T}\log{T}.$
Proof.
The Laurent expansion of $\zeta(s)$ at $s=1$ gives $(s-1)^{2}\zeta^{2}(s)=1+2\gamma(s-1)+(s-1)^{2}h(s)$ for some entire function $h(s)$, where $\gamma$ is the Euler–Mascheroni constant. Note that the preceding equality is valid for all $s\in\mathbb{C}$. Therefore,
$\operatorname*{Res}_{z=\frac{1}{2}-\mathrm{i}t}\left(\zeta^{2}{\left(\frac{1}{2}+\mathrm{i}t+z\right)}\Gamma(z)T^{z}\right)$ $=\lim_{z\to\frac{1}{2}-\mathrm{i}t}\frac{\mathrm{d}}{\mathrm{d}z}\left(\left(z-\frac{1}{2}+\mathrm{i}t\right)^{2}\zeta^{2}{\left(\frac{1}{2}+\mathrm{i}t+z\right)}\Gamma(z)T^{z}\right)$
$=\left(2\gamma+\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{2}-\mathrm{i}t\right)+\log{T}\right)T^{\frac{1}{2}-\mathrm{i}t}\Gamma\left(\frac{1}{2}-\mathrm{i}t\right).$
By [Olv74, Exercise 4.2 on p. 295] we have
$\left|\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{2}-\mathrm{i}t\right)\right|\leq\log{t}+\frac{\pi}{2}+\frac{1}{2}\log{\left(1+\frac{1}{4T^{2}}\right)}+\frac{1}{2T}+\frac{1}{3\sqrt{2}T^{2}}\leq\log{t}+1.7.$
By (72) we also have $\left|\Gamma\left(1/2-\mathrm{i}t\right)\right|\leq 3e^{-\frac{\pi}{2}t}$. The stated inequality now follows since $2\gamma+1.7+\log{T}\leq 3\log{t}$ and $e^{-\pi t/2}\log{t}$ is a decreasing function for $t\geq 5$. ∎
5.3.1. I. part
By Lemmas 1014 and 15 we have
$\left|2\int_{T}^{2T}\left|J_{1}(t)\right|^{2}\mathrm{d}t-\frac{1}{2\pi^{2}}T\log^{4}{T}\right|\leq\mathcal{J}_{1}(T_{0})T\log^{3}{T},$ (74)
where
$\mathcal{J}_{1}(T_{0})\mathrel{\mathop{:}}=2\eta(T_{0})\mathcal{D}_{6}(T_{0})+4\pi m_{0}\mathcal{D}_{4}(T_{0},0)+\frac{m_{0}\log{T_{0}}}{2\pi T_{0}},$ (75)
and $\mathcal{D}_{6}(x_{0})$ and $\mathcal{D}_{4}(x_{0},\sigma)$ are defined by (66) and (59), respectively. This shows that the main contribution in (12) comes from this part. In the subsequent sections, we show that all other terms are $\ll T\log^{3}{T}$.
5.3.2. II. part
Let $f(z)\mathrel{\mathop{:}}=\chi^{-1}(z)\sum_{n\leq T}n^{-z}d(n)$. It is clear that this is a holomorphic function in $\{z\in\mathbb{C}\colon|\Im\{z\}|>1\}$. By Cauchy’s theorem,
$\mathrm{i}\int_{T}^{2T}\left(J_{1}^{2}(t)+J_{2}^{2}(t)\right)\mathrm{d}t$ $=\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}+\int_{\frac{1}{2}-2\mathrm{i}T}^{\frac{1}{2}-\mathrm{i}T}\right)f^{2}(z)\mathrm{d}z$
$=\left(\int_{\sigma_{1}+\mathrm{i}T}^{\sigma_{1}+2\mathrm{i}T}+\int_{\sigma_{1}-2\mathrm{i}T}^{\sigma_{1}-\mathrm{i}T}\right)f^{2}(z)\mathrm{d}z$
$+\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{1}+\mathrm{i}T}+\int_{\frac{1}{2}-2\mathrm{i}T}^{\sigma_{1}-2\mathrm{i}T}\right)f^{2}(z)\mathrm{d}z+\left(\int_{\sigma_{1}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}+\int_{\sigma_{1}-\mathrm{i}T}^{\frac{1}{2}-\mathrm{i}T}\right)f^{2}(z)\mathrm{d}z.$ (76)
We have
$\left|\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{1}+\mathrm{i}T}+\int_{\frac{1}{2}-2\mathrm{i}T}^{\sigma_{1}-2\mathrm{i}T}\right)f^{2}(z)\mathrm{d}z\right|$ $\leq\int_{\sigma_{1}}^{\frac{1}{2}}\left(\left|\chi(u+\mathrm{i}T)\right|^{-2}+\left|\chi(u+2\mathrm{i}T)\right|^{-2}\right)\left(\sum_{n\leq T}n^{-u}d(n)\right)^{2}\mathrm{d}u$
$\leq\left(\frac{\mathcal{D}_{1}(T_{0},0)}{\mathcal{U}^{-}(T_{0})}\right)^{2}\left(\int_{\sigma_{1}}^{\frac{1}{2}}\frac{(2\pi)^{1-2u}+\pi^{1-2u}}{(1-u)^{2}}\mathrm{d}u\right)T\log^{2}{T}$
by Lemmas 11 and 18, with the same upper bound holding also for the modulus of the last pair of integrals in (76). Next,
$\left|\left(\int_{\sigma_{1}+\mathrm{i}T}^{\sigma_{1}+2\mathrm{i}T}+\int_{\sigma_{1}-2\mathrm{i}T}^{\sigma_{1}-\mathrm{i}T}\right)f^{2}(z)\mathrm{d}z\right|$ $\leq 2\int_{T}^{2T}\left|\chi(\sigma_{1}+\mathrm{i}u)\right|^{-2}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{1}+\mathrm{i}u}}\right|^{2}\mathrm{d}u$
$\leq\frac{2}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{T}{2\pi}\right)^{2\sigma_{1}-1}\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{1}+\mathrm{i}u}}\right|^{2}\mathrm{d}u$
$\leq\frac{2(2\pi)^{1-2\sigma_{1}}}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{1})+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{1}-1)\right)T\log^{3}{T}$
by Lemmas 1810 and 14. Therefore,
$\int_{T}^{2T}\left(J_{1}^{2}(t)+J_{2}^{2}(t)\right)\mathrm{d}t\leq\mathcal{J}_{2,1}(T_{0},\sigma_{1})T\log^{3}{T},$ (77)
where
$\mathcal{J}_{2,1}(T_{0},\sigma_{1})\mathrel{\mathop{:}}=\frac{2(2\pi)^{1-2\sigma_{1}}}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{1})+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{1}-1)\right)\\ +\frac{2}{\log{T_{0}}}\left(\frac{\mathcal{D}_{1}(T_{0},0)}{\mathcal{U}^{-}(T_{0})}\right)^{2}\int_{\sigma_{1}}^{\frac{1}{2}}\frac{(2\pi)^{1-2u}+\pi^{1-2u}}{(1-u)^{2}}\mathrm{d}u,$ (78)
and $\mathcal{D}_{4}(x_{0},\sigma)$, $\mathcal{D}_{1}(x_{0},\sigma_{0})$ and $\mathcal{U}^{-}(y_{0})$ are defined by (59), (55) and (73), respectively.
5.3.3. III. part
This part is the most challenging to estimate, and we need to divide it into several new terms. First, we estimate it for $k\in\{3,4\}$. Let
$g_{1}(z)\mathrel{\mathop{:}}=\chi(z)^{-1}\sum_{n>T}\frac{d(n)}{n^{z}}e^{-n/T},\quad g_{2}(z)\mathrel{\mathop{:}}=\chi(z)^{-1}\sum_{n\leq T}\frac{d(n)}{n^{z}}\left(e^{-n/T}-1\right).$
It is clear that $g_{1}(z)$ and $g_{2}(z)$ are holomorphic functions in $\{z\in\mathbb{C}\colon|\Im\{z\}|>1\}$. Define $F(z)\mathrel{\mathop{:}}=F_{1}(z)+F_{2}(z)$, where $F_{1}(z)\mathrel{\mathop{:}}=f(1-z)\left(g_{1}(z)+g_{2}(z)\right)$ and $F_{2}(z)\mathrel{\mathop{:}}=f(z)\left(g_{1}(z)+g_{2}(z)\right)$ with $f(z)$ defined as in Section 5.3.2. By Cauchy’s theorem,
$\mathrm{i}\int_{T}^{2T}\left(J_{1}(t)+J_{2}(t)\right)\left(J_{3}(t)+J_{4}(t)\right)\mathrm{d}t=\int_{\frac{1}{2}+\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}F(z)\mathrm{d}z=\int_{\sigma_{2}+\mathrm{i}T}^{\sigma_{2}+2\mathrm{i}T}F_{1}(z)\mathrm{d}z+\int_{\sigma_{3}+\mathrm{i}T}^{\sigma_{3}+2\mathrm{i}T}F_{2}(z)\mathrm{d}z\\ +\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{2}+\mathrm{i}T}+\int_{\sigma_{2}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)F_{1}(z)\mathrm{d}z+\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{3}+\mathrm{i}T}+\int_{\sigma_{3}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)F_{2}(z)\mathrm{d}z.$ (79)
We need to estimate the moduli of each of the above terms. Starting with the first two terms on the right-hand side of (79), we have
$\left|\int_{\sigma_{2}+\mathrm{i}T}^{\sigma_{2}+2\mathrm{i}T}F_{1}(z)\mathrm{d}z\right|\leq\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(2+\frac{1}{T_{0}}\right)^{\sigma_{2}-\frac{1}{2}}\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{1-\sigma_{2}+\mathrm{i}u}}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\times\\ \times\left(\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{\sigma_{2}+\mathrm{i}u}}e^{-n/T}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}+\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{2}+\mathrm{i}u}}\left(e^{-n/T}-1\right)\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\right)$
and
$\left|\int_{\sigma_{3}+\mathrm{i}T}^{\sigma_{3}+2\mathrm{i}T}F_{2}(z)\mathrm{d}z\right|\leq\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{2\pi}{T}\right)^{1-2\sigma_{3}}\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\times\\ \times\left(\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}e^{-n/T}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}+\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}\left(e^{-n/T}-1\right)\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\right).$
by the Cauchy–Bunyakovsky–Schwarz (CBS) inequality and Lemma 18. Then
$\left|\int_{1+\mathrm{i}T}^{1+2\mathrm{i}T}F_{1}(z)\mathrm{d}z+\int_{\mathrm{i}T}^{2\mathrm{i}T}F_{2}(z)\mathrm{d}z\right|\leq\mathfrak{G}_{1}\left(T_{0},\bm{\sigma^{\prime}}\right)T\log^{3}{T}$ (80)
by Lemmas 101416 and 17, where
$\mathfrak{G}_{1}\left(T_{0},\bm{\sigma^{\prime}}\right)$ $\mathrel{\mathop{:}}=\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(2+\frac{1}{T_{0}}\right)^{\sigma_{2}-\frac{1}{2}}\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2-2\sigma_{2})+2\pi m_{0}\mathcal{D}_{4}(T_{0},1-2\sigma_{2})\right)^{\frac{1}{2}}\times$
×$\times\left(\left(\eta(T_{0})+2\pi m_{0}\right)^{\frac{1}{2}}\mathcal{D}_{7}(T_{0})^{\frac{1}{2}}+\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{2}-2)+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{2}-3)\right)^{\frac{1}{2}}\right)$
$+\frac{(2\pi)^{1-2\sigma_{3}}}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{3})+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{3}-1)\right)^{\frac{1}{2}}\times$
×$\times\left(\left(\eta(T_{0})+2\pi m_{0}\right)^{\frac{1}{2}}\mathcal{D}_{7}(T_{0})^{\frac{1}{2}}+\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{3}-2)+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{3}-3)\right)^{\frac{1}{2}}\right).$ (81)
Here, $\mathcal{D}_{4}(x_{0},\sigma)$, $\mathcal{D}_{7}(x_{0})$ and $\mathcal{U}^{-}(y_{0})$ are defined by (59), (67) and (73), respectively. Turning to the last two terms in (79), we have
$\left|\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{2}+\mathrm{i}T}+\int_{\sigma_{2}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)F_{1}(z)\mathrm{d}z\right|\leq\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\int_{\frac{1}{2}}^{\sigma_{2}}\left(\left(1+\frac{1}{T_{0}}\right)^{u-\frac{1}{2}}+\left(1+\frac{1}{2T_{0}}\right)^{u-\frac{1}{2}}\right)\times\\ \times\left(\sum_{n>T}\frac{d(n)}{n^{u}}e^{-n/T}+\sum_{n\leq T}\frac{d(n)}{n^{u}}\left(1-e^{-n/T}\right)\right)\sum_{n\leq T}\frac{d(n)}{n^{1-u}}\mathrm{d}u$
and
$\left|\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{3}+\mathrm{i}T}+\int_{\sigma_{3}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)F_{2}(z)\mathrm{d}z\right|\leq\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\int_{\sigma_{3}}^{\frac{1}{2}}\left((2\pi)^{1-2u}+\pi^{1-2u}\right)T^{2u-1}\times\\ \times\left(\sum_{n>T}\frac{d(n)}{n^{u}}e^{-n/T}+\sum_{n\leq T}\frac{d(n)}{n^{u}}\left(1-e^{-n/T}\right)\right)\sum_{n\leq T}\frac{d(n)}{n^{u}}\mathrm{d}u$
by Lemma 18. Then
$\left|\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{2}+\mathrm{i}T}+\int_{\sigma_{2}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)F_{1}(z)\mathrm{d}z+\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{3}+\mathrm{i}T}+\int_{\sigma_{3}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)F_{2}(z)\mathrm{d}z\right|\leq\mathfrak{G}_{2}\left(T_{0},\bm{\sigma^{\prime}}\right)T\log^{2}{T}$ (82)
by Lemmas 11 and 12, where
$\mathfrak{G}_{2}\left(T_{0},\bm{\sigma^{\prime}}\right)\mathrel{\mathop{:}}=\frac{\mathcal{D}_{1}(T_{0},0)}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\int_{\frac{1}{2}}^{\sigma_{2}}\left(\left(1+\frac{1}{T_{0}}\right)^{u-\frac{1}{2}}+\left(1+\frac{1}{2T_{0}}\right)^{u-\frac{1}{2}}\right)\left(\frac{\mathcal{D}_{1}(T_{0},u-1)}{2-u}+\mathcal{D}_{3}(T_{0})\right)\frac{\mathrm{d}u}{u}\\ +\frac{\mathcal{D}_{1}(T_{0},0)}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\int_{\sigma_{3}}^{\frac{1}{2}}\left((2\pi)^{1-2u}+\pi^{1-2u}\right)\left(\frac{\mathcal{D}_{1}(T_{0},u-1)}{2-u}+\mathcal{D}_{3}(T_{0})\right)\frac{\mathrm{d}u}{1-u}.$ (83)
Here, $\mathcal{D}_{1}(x_{0},\sigma_{0})$, $\mathcal{D}_{3}(x_{0})$ and $\mathcal{U}^{-}(y_{0})$ are defined by (55), (58) and (73), respectively.
It remains to estimate the third part also for $k\in\{5,6\}$, and the procedure is very similar to the one in the preceding section. Define
$g_{3}(z)\mathrel{\mathop{:}}=-\frac{1}{2\pi}\chi^{-1}(z)\int_{-\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}}^{\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}}\chi^{2}(z+c_{1}+\mathrm{i}u)\Gamma(c_{1}+\mathrm{i}u)T^{c_{1}+\mathrm{i}u}\sum_{n>T}\frac{d(n)}{n^{1-z-c_{1}-\mathrm{i}u}}\mathrm{d}u,$
$g_{4}(z)\mathrel{\mathop{:}}=-\frac{1}{2\pi}\chi^{-1}(z)\int_{-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}}^{\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}}\chi^{2}(z+c_{2}+\mathrm{i}u)\Gamma(c_{2}+\mathrm{i}u)T^{c_{2}+\mathrm{i}u}\sum_{n\leq T}\frac{d(n)}{n^{1-z-c_{2}-\mathrm{i}u}}\mathrm{d}u.$
It is clear that $g_{3}(z)$ and $g_{4}(z)$ are holomorphic functions in $\{z\in\mathbb{C}\colon 1/2-c_{2}<\Re\{z\}<|c_{1}|,\Im\{z\}>1+\mathfrak{b_{1}}(\log{T})^{\mathfrak{b}_{2}}\}$. Define $G(z)\mathrel{\mathop{:}}=G_{1}(z)+G_{2}(z)$, where $G_{1}(z)\mathrel{\mathop{:}}=f(1-z)\left(g_{3}(z)+g_{4}(z)\right)$ and $G_{2}(z)\mathrel{\mathop{:}}=f(z)\left(g_{3}(z)+g_{4}(z)\right)$ with $f(z)$ defined as in Section 5.3.2. By Cauchy’s theorem,
$\mathrm{i}\int_{T}^{2T}\left(J_{1}(t)+J_{2}(t)\right)\left(J_{5}(t)+J_{6}(t)\right)\mathrm{d}t=\int_{\frac{1}{2}+\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}G(z)\mathrm{d}z=\int_{\sigma_{4}+\mathrm{i}T}^{\sigma_{4}+2\mathrm{i}T}G_{1}(z)\mathrm{d}z+\int_{\sigma_{5}+\mathrm{i}T}^{\sigma_{5}+2\mathrm{i}T}G_{2}(z)\mathrm{d}z\\ +\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{4}+\mathrm{i}T}+\int_{\sigma_{4}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)G_{1}(z)\mathrm{d}z+\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{5}+\mathrm{i}T}+\int_{\sigma_{5}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)G_{2}(z)\mathrm{d}z.$ (84)
We need to estimate the moduli of each of the above terms. Starting with the first term on the right-hand side of (84), note that $20<1-\sigma_{4}<1/$ 2. By the CBS inequality and Lemma 18 we then have
$\left|\int_{\sigma_{4}+\mathrm{i}T}^{\sigma_{4}+2\mathrm{i}T}G_{1}(z)\mathrm{d}z\right|\leq\frac{1}{\mathcal{U}^{-}(T_{0})}\left(\frac{2\pi}{T}\right)^{\sigma_{4}-\frac{1}{2}}\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{1-\sigma_{4}-\mathrm{i}t}}\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\times\\ \times\left(\left(\int_{T}^{2T}\left|g_{3}(\sigma_{4}+\mathrm{i}t)\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}+\left(\int_{T}^{2T}\left|g_{4}(\sigma_{4}+\mathrm{i}t)\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\right).$ (85)
We estimate the second integral on the right-hand side of (85). Note that $-1/2<c_{1}+\sigma_{4}<0$, implying $|c_{1}+\sigma_{4}+\mathrm{i}(t+u)|\leq 1/2+2T+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ since $t\in[T,2T]$ and $|u|\leq\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}$. By the CBS inequality and Lemma 18 we thus have
$\int_{T}^{2T}\left|g_{3}(\sigma_{4}+\mathrm{i}t)\right|^{2}\mathrm{d}t$ $\leq\frac{T^{1-2\sigma_{4}+2|c_{1}|}}{\left(2\pi\cdot\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{1}{\pi}+\frac{1}{2\pi T_{0}}\right)^{2\sigma_{4}-1}\left(\frac{1}{\pi}+\frac{\frac{1}{2}+\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{2\left(1-2\sigma_{4}\right)+4|c_{1}|}\times$
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\times$
×$\times\int_{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{1-\sigma_{4}-c_{1}-\mathrm{i}(t+u)}}\right|^{2}\mathrm{d}t\right)\mathrm{d}u.$
Moreover,
$\int_{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{1-\sigma_{4}-c_{1}-\mathrm{i}(t+u)}}\right|^{2}\mathrm{d}t\right)\mathrm{d}u\leq\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\times\\ \times\left(\frac{\eta(T_{0})\mathcal{D}_{5}\left(T_{0},2-2\sigma_{4}+2|c_{1}|\right)}{(1-2\sigma_{4}+2|c_{1}|)^{4}}+\frac{2\pi m_{0}\mathcal{D}_{5}\left(T_{0},1-2\sigma_{4}+2|c_{1}|\right)}{\left(2|c_{1}|-2\sigma_{4}\right)^{4}}\right)T^{2\sigma_{4}-2|c_{1}|}\log^{3}{T}$
by Lemmas 10 and 14. Therefore,
$\int_{T}^{2T}\left|g_{3}(\sigma_{4}+\mathrm{i}t)\right|^{2}\mathrm{d}t\leq\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)T\log^{3}{T},$ (86)
where
$\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)$ $\mathrel{\mathop{:}}=\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{1}{\pi}+\frac{1}{2\pi T_{0}}\right)^{2\sigma_{4}-1}\left(\frac{1}{\pi}+\frac{\frac{1}{2}+\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{2\left(1-2\sigma_{4}\right)+4|c_{1}|}\times$
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\frac{\eta(T_{0})\mathcal{D}_{5}\left(T_{0},2-2\sigma_{4}+2|c_{1}|\right)}{(2\pi)^{2}(1-2\sigma_{4}+2|c_{1}|)^{4}}+\frac{m_{0}\mathcal{D}_{5}\left(T_{0},1-2\sigma_{4}+2|c_{1}|\right)}{2\pi\left(2|c_{1}|-2\sigma_{4}\right)^{4}}\right).$ (87)
Here, $\mathcal{D}_{5}(x_{0},\sigma)$ and $\mathcal{U}^{\pm}(y_{0})$ are defined by (14) and (73), respectively. Turning to the third integral on the right-hand side of (85), note that 1/$21/2<c_{2}+\sigma_{4}<3/$ 2 and $|\sigma_{4}+c_{2}+\mathrm{i}(t+u)|\geq T-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ since $t\in[T,2T]$ and $|u|\leq\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}$. Also, $-1<2(1-\sigma_{4}-c_{2})<1$. Similarly as before, the CBS inequality and Lemmas 1014 and 18 imply
$\int_{T}^{2T}\left|g_{4}(\sigma_{4}+\mathrm{i}t)\right|^{2}\mathrm{d}t\leq\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)T\log^{3}{T},$ (88)
where
$\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)$ $\mathrel{\mathop{:}}=\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{1}{\pi}+\frac{1}{2\pi T_{0}}\right)^{2\sigma_{4}-1}\left(\frac{1}{2\pi}-\frac{\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}}{2\pi T_{0}}\right)^{2(1-2(\sigma_{4}+c_{2}))}\times$
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\frac{\eta(T_{0})\mathcal{D}_{4}\left(T_{0},2\left(1-\sigma_{4}-c_{2}\right)\right)}{(2\pi)^{2}}+\frac{m_{0}\mathcal{D}_{4}\left(T_{0},1-2\left(\sigma_{4}+c_{2}\right)\right)}{2\pi}\right).$ (89)
Here, $\mathcal{D}_{4}(x_{0},\sigma)$ and $\mathcal{U}^{\pm}(y_{0})$ are defined by (59) and (73), respectively. Note that (86) and (88) also hold for $\sigma_{4}=1/2$. This fact will be used in Section 5.3.4. Using (86) and (88) in (85), together with Lemmas 10 and 14 to estimate the first integral on the right-hand side of (85), gives
$\left|\int_{\sigma_{4}+\mathrm{i}T}^{\sigma_{4}+2\mathrm{i}T}G_{1}(z)\mathrm{d}z\right|\leq\mathfrak{G}_{3}\left(T_{0},\sigma_{4},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{T},$ (90)
where
$\mathfrak{G}_{3}\left(T_{0},\sigma_{4},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ $\mathrel{\mathop{:}}=\frac{(2\pi)^{\sigma_{4}}}{\mathcal{U}^{-}(T_{0})}\left(\frac{\eta(T_{0})}{2\pi}\mathcal{D}_{4}(T_{0},2(1-\sigma_{4}))+m_{0}\mathcal{D}_{4}(T_{0},1-2\sigma_{4})\right)^{\frac{1}{2}}\times$
×$\times\left(\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)^{\frac{1}{2}}+\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)^{\frac{1}{2}}\right).$ (91)
Here, $\mathcal{D}_{4}(x_{0},\sigma)$, $\mathcal{U}^{-}(y_{0})$, $\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (59), (73), (5.3.3) and (5.3.3), respectively. Moving to the second term on the right-hand side of (84), we have
$\left|\int_{\sigma_{5}+\mathrm{i}T}^{\sigma_{5}+2\mathrm{i}T}G_{2}(z)\mathrm{d}z\right|\leq\frac{1}{\mathcal{U}^{-}(T_{0})}\left(\frac{2\pi}{T}\right)^{\frac{1}{2}-\sigma_{5}}\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{5}+\mathrm{i}t}}\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\times\\ \times\left(\left(\int_{T}^{2T}\left|g_{3}(\sigma_{5}+\mathrm{i}t)\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}+\left(\int_{T}^{2T}\left|g_{4}(\sigma_{5}+\mathrm{i}t)\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\right)$
by the CBS inequality and Lemma 18 since $20<\sigma_{5}<1/$ 2. Estimation of the second and the third integral is similar as before. We obtain
$\int_{T}^{2T}\left|g_{3}(\sigma_{5}+\mathrm{i}t)\right|^{2}\mathrm{d}t\leq\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)T\log^{3}{T}$
and
$\int_{T}^{2T}\left|g_{4}(\sigma_{5}+\mathrm{i}t)\right|^{2}\mathrm{d}t\leq\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)T\log^{3}{T}$
by Lemmas 1014 and 18, where
$\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)$ $\mathrel{\mathop{:}}=\frac{1}{(2\pi)^{1+2\sigma_{5}}\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{1}{\pi}+\frac{1+\mathfrak{a}_{1}\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{2\left(1-2\sigma_{5}\right)+4|c_{1}|}\times$
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\frac{\eta(T_{0})\mathcal{D}_{5}(T_{0},2-2\sigma_{5}+2|c_{1}|)}{\left(1-2\sigma_{5}+2|c_{1}|\right)^{4}}+\frac{2\pi m_{0}\mathcal{D}_{5}(T_{0},1-2\sigma_{5}+2|c_{1}|)}{\left(2|c_{1}|-2\sigma_{5}\right)^{4}}\right)$ (92)
and
$\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)$ $\mathrel{\mathop{:}}=\frac{1}{(2\pi)^{1+2\sigma_{5}}\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{1}{2\pi}-\frac{\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}}{2\pi T_{0}}\right)^{2(1-2(\sigma_{5}+c_{2}))}\times$
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\eta(T_{0})\mathcal{D}_{4}\left(T_{0},2\left(1-\sigma_{5}-c_{2}\right)\right)+2\pi m_{0}\mathcal{D}_{4}\left(T_{0},1-2\left(\sigma_{5}+c_{2}\right)\right)\right).$ (93)
Here, $\mathcal{D}_{5}(x_{0},\sigma)$, $\mathcal{D}_{4}(x_{0},\sigma)$ and $\mathcal{U}^{\pm}(y_{0})$ are defined by (14), (59) and (73), respectively. Therefore,
$\left|\int_{\sigma_{5}+\mathrm{i}T}^{\sigma_{5}+2\mathrm{i}T}G_{2}(z)\mathrm{d}z\right|\leq\mathfrak{G}_{4}\left(T_{0},\sigma_{5},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{T},$ (94)
where
$\mathfrak{G}_{4}\left(T_{0},\sigma_{5},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)$ $\mathrel{\mathop{:}}=\frac{(2\pi)^{\frac{1}{2}-\sigma_{5}}}{\mathcal{U}^{-}(T_{0})}\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{5})+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{5}-1)\right)^{\frac{1}{2}}\times$
×$\times\left(\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)^{\frac{1}{2}}+\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)^{\frac{1}{2}}\right).$ (95)
Here, $\mathcal{D}_{4}(x_{0},\sigma)$, $\mathcal{U}^{-}(y_{0})$, $\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (59), (73), (5.3.3) and (5.3.3), respectively. It remains to estimate the last two terms on the right-hand side of (84). Beginning with the first such term, note that by writing $z=\sigma+\mathrm{i}T$ in the case of the first integral and $z=\sigma+2\mathrm{i}T$ in the case of the second integral, we have $\sigma\in[1/2,\sigma_{4}]$, implying $-1/2<\sigma+c_{1}<0$, 1/$21/2<\sigma+c_{2}<3/$ 2, $|\sigma+c_{1}+\mathrm{i}(u+kT)|\leq 1/2+kT+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$, and $|\sigma+c_{2}+\mathrm{i}(u+kT)|\geq kT-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$. Define
$\mathfrak{g}_{5}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)\mathrel{\mathop{:}}=\int_{\frac{1}{2}}^{\sigma_{4}}\left(\left(1+\frac{1}{T_{0}}\right)^{\sigma-\frac{1}{2}}\left(\frac{1}{2\pi}+\frac{\frac{1}{2}+\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma-|c_{1}|)}\right.\\ \left.+\left(1+\frac{1}{2T_{0}}\right)^{\sigma-\frac{1}{2}}\left(\frac{1}{\pi}+\frac{\frac{1}{2}+\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma-|c_{1}|)}\right)\frac{\mathcal{D}_{2}(T_{0},1-\sigma+|c_{1}|)}{\sigma\left(\sigma-|c_{1}|\right)^{2}}\mathrm{d}\sigma,$ (96)
where $\mathcal{D}_{2}(x_{0},\sigma)$ is defined by (56), and
$\mathfrak{g}_{6}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\int_{\frac{1}{2}}^{\sigma_{4}}\left(\left(1+\frac{1}{T_{0}}\right)^{\sigma-\frac{1}{2}}\left(\frac{1}{2\pi}-\frac{\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma+c_{2})}\right.\\ \left.+\left(1+\frac{1}{2T_{0}}\right)^{\sigma-\frac{1}{2}}\left(\frac{1}{\pi}-\frac{\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma+c_{2})}\right)\frac{\mathrm{d}\sigma}{\sigma(\sigma+c_{2})}.$ (97)
By Lemmas 11 and 18 we thus have
$\left|\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{4}+\mathrm{i}T}+\int_{\sigma_{4}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)G_{1}(z)\mathrm{d}z\right|\leq\mathfrak{G}_{5}\left(T_{0},\sigma_{4},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{2}{T},$ (98)
where
$\mathfrak{G}_{5}\left(T_{0},\sigma_{4},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\frac{\mathcal{D}_{1}(T_{0},0)}{2\pi}\left(\frac{\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)}{\mathcal{U}^{-}(T_{0})}\right)^{2}\mathfrak{g}_{5}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\\ +\frac{\mathcal{D}_{1}(T_{0},0)\mathcal{D}_{1}(T_{0},-1/2)}{2\pi}\left(\frac{\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}}\right)}{\mathcal{U}^{-}(T_{0})}\right)^{2}\mathfrak{g}_{6}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u.$ (99)
Here, $\mathcal{D}_{1}(x_{0},\sigma_{0})$, $\mathcal{U}^{\pm}(y_{0})$, $\mathfrak{g}_{5}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{6}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (55), (73), (96) and (97), respectively. Turning to the second term on the right-hand side of (84), we have, similarly as before, that $\sigma\in[\sigma_{5},1/2]$, implying $-1<\sigma+c_{1}<0$, $11/2<\sigma+c_{2}<$ 1, $|\sigma+c_{1}+\mathrm{i}(u+kT)|\leq 1+kT+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$, and $|\sigma+c_{2}+\mathrm{i}(u+kT)|\geq kT-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$. Define
$\mathfrak{g}_{7}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)\mathrel{\mathop{:}}=\int_{\sigma_{5}}^{\frac{1}{2}}\left((2\pi)^{1-2\sigma}\left(\frac{1}{2\pi}+\frac{1+\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma-|c_{1}|)}\right.\\ \left.+\pi^{1-2\sigma}\left(\frac{1}{\pi}+\frac{1+\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma-|c_{1}|)}\right)\frac{\mathcal{D}_{2}(T_{0},1-\sigma+|c_{1}|)}{(1-\sigma)\left(\sigma-|c_{1}|\right)^{2}}\mathrm{d}\sigma,$ (100)
where $\mathcal{D}_{2}(x_{0},\sigma)$ is defined by (56), and
$\mathfrak{g}_{8}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\int_{\sigma_{5}}^{\frac{1}{2}}\left((2\pi)^{1-2\sigma}\left(\frac{1}{2\pi}-\frac{\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma+c_{2})}\right.\\ \left.+\pi^{1-2\sigma}\left(\frac{1}{\pi}-\frac{\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}}}{2\pi T_{0}}\right)^{1-2(\sigma+c_{2})}\right)\frac{\mathrm{d}\sigma}{(1-\sigma)(\sigma+c_{2})}.$ (101)
By Lemmas 11 and 18 we thus have
$\left|\left(\int_{\frac{1}{2}+\mathrm{i}T}^{\sigma_{5}+\mathrm{i}T}+\int_{\sigma_{5}+2\mathrm{i}T}^{\frac{1}{2}+2\mathrm{i}T}\right)G_{2}(z)\mathrm{d}z\right|\leq\mathfrak{G}_{6}\left(T_{0},\sigma_{5},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{2}{T},$ (102)
where
$\mathfrak{G}_{6}\left(T_{0},\sigma_{5},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\frac{\mathcal{D}_{1}(T_{0},0)}{2\pi}\left(\frac{\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)}{\mathcal{U}^{-}(T_{0})}\right)^{2}\mathfrak{g}_{7}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\\ +\frac{1}{2\pi}\left(\frac{\mathcal{D}_{1}(T_{0},0)\cdot\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}}\right)}{\mathcal{U}^{-}(T_{0})}\right)^{2}\mathfrak{g}_{8}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u.$ (103)
Here, $\mathcal{D}_{1}(x_{0},\sigma_{0})$, $\mathcal{U}^{\pm}(y_{0})$, $\mathfrak{g}_{7}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{8}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (55), (73), (100) and (101), respectively.
Finally, it remains to collect the relevant estimates in order to bound the left-hand sides of (79) and (84). For the former equality, we are using (80) and (82), while for the latter equality we are using (90), (94), (98) and (102). The final result is that
$2\sum_{k=3}^{6}\left|\int_{T}^{2T}\left(J_{1}(t)+J_{2}(t)\right)J_{k}(t)\mathrm{d}t\right|\leq 2\mathcal{J}_{2,2}\left(T_{0},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{T},$ (104)
where
$\mathcal{J}_{2,2}\left(T_{0},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)=\mathfrak{G}_{1}\left(T_{0},\bm{\sigma}^{\prime}\right)+\mathfrak{G}_{3}\left(T_{0},\sigma_{4},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)+\mathfrak{G}_{4}\left(T_{0},\sigma_{5},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\\ +\frac{1}{\log{T_{0}}}\left(\mathfrak{G}_{2}\left(T_{0},\bm{\sigma}^{\prime}\right)+\mathfrak{G}_{5}\left(T_{0},\sigma_{4},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)+\mathfrak{G}_{6}\left(T_{0},\sigma_{5},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\right).$ (105)
Here, the functions $\mathfrak{G}_{1},\ldots,\mathfrak{G}_{6}$ are defined by (5.3.3), (83), (5.3.3), (5.3.3), (99) and (103), respectively.
5.3.4. IV. part
By Lemmas 1016 and 17 we have
$\frac{1}{T\log^{3}{T}}\int_{T}^{2T}\left|J_{3}(t)\right|^{2}\mathrm{d}t\leq\mathcal{J}_{3}(T_{0})\mathrel{\mathop{:}}=\left(\eta(T_{0})+2\pi m_{0}\right)\mathcal{D}_{7}(T_{0}),$ (106)
where $\mathcal{D}_{7}(x_{0})$ is defined by (67), and
$\frac{1}{T\log^{3}{T}}\int_{T}^{2T}\left|J_{4}(t)\right|^{2}\mathrm{d}t\leq\mathcal{J}_{4}(T_{0})\mathrel{\mathop{:}}=\eta(T_{0})\mathcal{D}_{4}(T_{0},-1)+2\pi m_{0}\mathcal{D}_{4}(T_{0},-2),$ (107)
where $\mathcal{D}_{4}(x_{0},\sigma)$ is defined by (59). Turning to the integral of $|J_{5}(t)|^{2}$, observe that
$\frac{1}{T\log^{3}{T}}\int_{T}^{2T}\left|J_{5}(t)\right|^{2}\mathrm{d}t=\frac{1}{T\log^{3}{T}}\int_{T}^{2T}\left|g_{3}\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t\leq\mathcal{J}_{5}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)\mathrel{\mathop{:}}=\mathfrak{g}_{1}\left(T_{0},\frac{1}{2},c_{1},\bm{\mathfrak{a}}\right)$ (108)
by (86), where $\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)$ is defined by (5.3.3). Similarly,
$\frac{1}{T\log^{3}{T}}\int_{T}^{2T}|J_{6}(t)|^{2}\mathrm{d}t=\frac{1}{T\log^{3}{T}}\int_{T}^{2T}\left|g_{4}\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t\leq\mathcal{J}_{6}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\mathfrak{g}_{2}\left(T_{0},\frac{1}{2},c_{2},\bm{\mathfrak{b}}\right)$ (109)
by (88), where $\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)$ is defined by (5.3.3).
5.3.5. V. part
Using estimates (106)–(109), we obtain
$2\mathop{\sum\sum}_{3\leq n<m\leq 6}\int_{T}^{2T}\left|J_{n}(t)J_{m}(t)\right|\mathrm{d}t\leq\left(2\mathop{\sum\sum}_{3\leq n<m\leq 6}\left(\mathcal{J}_{n}\mathcal{J}_{m}\right)^{\frac{1}{2}}\right)T\log^{3}{T}$ (110)
by the CBS inequality, where we borrowed the shorter notation for $\mathcal{J}_{n}$ from Theorem 2.
5.3.6. VI. part
By Lemma 18 we have
$\left|\chi\left(\frac{1}{2}+c_{1}+\mathrm{i}(t+u)\right)\right|\leq(2T)^{|c_{1}|}\left(0.4+\frac{16}{\left(2T_{0}+\mathfrak{a}_{1}\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}}\right)^{|c_{1}|}}\right)\left(1+\frac{|u|}{2T_{0}}\right)^{|c_{1}|}$
and
$\left|\Gamma(c_{1}+\mathrm{i}u)\right|\leq\sqrt{2\pi}\exp{\left(\frac{1}{2\mathfrak{a}_{1}^{2}\left(\log{T_{0}}\right)^{2\mathfrak{a}_{2}}}\right)}T^{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}-1}}\frac{e^{-|u|/2}}{|u|^{|c_{1}|+1/2}}$
since $c_{1}\in(-1,-1/2)$, $t\in[T,2T]$ and $|u|\geq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$. Therefore,
$\frac{1}{2\pi}\left|\left(\int_{-\infty}^{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}+\int_{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\infty}\right)P(0,1,c_{1};u)\mathrm{d}u\right|\leq\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)T^{\frac{1}{2}-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}-1}}\log{T}$ (111)
by Lemma 11, where
$\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)\mathrel{\mathop{:}}=2^{2|c_{1}|}\sqrt{\frac{2}{\pi}}\left(0.4+\frac{16}{\left(2T_{0}+\mathfrak{a}_{1}\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}}\right)^{|c_{1}|}}\right)^{2}\exp{\left(\frac{1}{2\mathfrak{a}_{1}^{2}\left(\log{T_{0}}\right)^{2\mathfrak{a}_{2}}}\right)}\times\\ \times\frac{\mathcal{D}_{2}\left(T_{0},\frac{1}{2}+|c_{1}|\right)}{\left(|c_{1}|-\frac{1}{2}\right)^{2}}\int_{4}^{\infty}\left(1+\frac{u}{2T_{0}}\right)^{2|c_{1}|}\frac{e^{-\frac{1}{2}u}}{u^{|c_{1}|+\frac{1}{2}}}\mathrm{d}u$ (112)
since $\mathfrak{a}_{1}\left(\log{T_{0}}\right)^{\mathfrak{a}_{2}}\geq\log{T_{0}}\geq 4$. Here, $\mathcal{D}_{2}(x_{0},\sigma_{0})$ is defined by (56). By Lemma 18 we also have
$\left|\chi\left(\frac{1}{2}+c_{1}+\mathrm{i}(t+u)\right)\right|\leq\frac{6.31}{\frac{1}{2}-c_{2}}$
and
$\left|\Gamma(c_{2}+\mathrm{i}u)\right|\leq\sqrt{2\pi}\exp{\left(\frac{1}{2\mathfrak{b}_{1}^{2}\left(\log{T_{0}}\right)^{2\mathfrak{b}_{2}}}\right)}\left(\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)^{c_{2}-\frac{1}{2}}T^{-\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}-1}}e^{-\frac{1}{2}|u|}$
since $c_{2}\in(0,1/2)$, $t\in[T,2T]$ and $|u|\geq\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}$. Therefore,
$\frac{1}{2\pi}\left|\left(\int_{-\infty}^{-\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}}+\int_{\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}}^{\infty}\right)P(1,0,c_{2};u)\mathrm{d}u\right|\leq\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)T^{\frac{1}{2}+2c_{2}-\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}-1}}\log{T}$ (113)
by Lemma 11, where
$\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\frac{2}{e^{4}}\sqrt{\frac{2}{\pi}}\left(\frac{6.31}{\frac{1}{2}-c_{2}}\right)^{2}\exp{\left(\frac{1}{2\mathfrak{b}_{1}^{2}\left(\log{T_{0}}\right)^{2\mathfrak{b}_{2}}}\right)}\left(\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)^{c_{2}-\frac{1}{2}}\frac{\mathcal{D}_{1}(T_{0},0)}{\frac{1}{2}+c_{2}}$ (114)
since $\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\geq 2\log{T_{0}}\geq 8$. Here, $\mathcal{D}_{1}(x_{0},\sigma_{0})$ is defined by (55). Then we have
$2\int_{T}^{2T}\left|J_{0}(t)\right|\cdot\left|\zeta\left(\frac{1}{2}+\mathrm{i}t\right)\right|^{2}\mathrm{d}t+\int_{T}^{2T}\left|J_{0}(t)\right|^{2}\mathrm{d}t\leq\mathcal{J}_{0}\left(T_{0},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)T\log^{3}{T},$ (115)
where
$\mathcal{J}_{0}\left(T_{0},\bm{c},\bm{\mathfrak{a}},\bm{\mathfrak{b}}\right)\mathrel{\mathop{:}}=\frac{7}{2}\left(\left(1+\frac{3}{14}\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)\right)\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)T_{0}^{1-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}-1}}\right.\\ \left.+\left(1+\frac{3}{14}\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)\right)\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)T_{0}^{2-\mathfrak{b}_{1}(\log{T_{0}})^{\mathfrak{b}_{2}-1}}+12.1T_{0}e^{-\frac{\pi}{2}T_{0}}\right),$ (116)
and $\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (112) and (114), respectively. In the derivation of (115), we have used inequalities (111) and (113), together with Lemma 19 and $\left|\zeta(1/2+\mathrm{i}t)\right|\leq(2T)^{1/6}\log{(2T)}$ for $t\in[T,2T]$, see [HPY24, Theorem 1.1].
5.3.7. Proof of estimate (12)
First, by (74) we have
$\left|\mathcal{M}_{2}(T,2T)-\frac{1}{2\pi^{2}}T\log^{4}{T}\right|\leq\mathcal{J}_{1}(T_{0})T\log^{3}{T}+\left|\mathcal{M}_{2}(T,2T)-2\int_{T}^{2T}\left|J_{1}(t)\right|^{2}\mathrm{d}t\right|,$
where $\mathcal{J}_{1}(T_{0})$ is defined by (75). We are using (52) and (53) in order to estimate the last term in the last inequality. Then (12) follows by (77), (104), (106), (107), (108), (109), (110) and (115).
5.3.8. Proof of estimate (13)
We are using (54). By (74) we have
$4\int_{T}^{2T}\left|J_{1}(t)\right|^{2}\mathrm{d}t\leq\frac{1}{\pi^{2}}T\log^{4}{T}+2\mathcal{J}_{1}(T_{0})T\log^{3}{T}.$
Also, the above inequality, together with the CBS inequality and the inequalities from Section 5.3.4, implies
$4\sum_{k=3}^{6}\int_{T}^{2T}\left|J_{1}(t)J_{k}(t)\right|\mathrm{d}t$ $\leq 2\left(4\int_{T}^{2T}\left|J_{1}(t)\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\sum_{k=3}^{6}\left(\int_{T}^{2T}\left|J_{k}(t)\right|^{2}\mathrm{d}t\right)^{\frac{1}{2}}$
$\leq 2\left(T\log^{3}{T}\right)\sum_{k=3}^{6}\left(\frac{\mathcal{J}_{k}}{\pi^{2}}\log{T}+2\mathcal{J}_{1}\mathcal{J}_{k}\right)^{\frac{1}{2}}.$
Other terms in (54) are estimated as in the proof of (12), that is, we are using estimates from Sections 5.3.45.3.5 and 5.3.6. The proof is thus complete.